### v1

# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

Abstract
Recent advancements in large language models (LLMs) have shown strong general reasoning abilities, yet their effectiveness in financial reasoning remains underexplored. In this study, we comprehensively evaluate 16 powerful reasoning and general LLMs on three complex financial tasks involving financial text, tabular data, and equations, assessing numerical reasoning, tabular interpretation, financial terminology comprehension, long-context processing, and equation-based problem solving. Our results show that while better datasets and pretraining improve financial reasoning, general enhancements like CoT fine-tuning do not always yield consistent gains. Moreover, all reasoning strategies face challenges in improving performance on long-context and multi-table tasks. To address these limitations, we develop a financial reasoning-enhanced model based on Llama-3.1-8B-Instruct, by CoT fine-tuning and reinforcement learning with domain-specific reasoning paths. Even with simple fine-tuning with one financial dataset, our model achieves a consistent 10% performance improvement across tasks, surpassing all 8B models and even Llama3-70B-Instruct and Llama3.1-70B-Instruct on average. Our results highlight the need for domain-specific adaptations in financial tasks, emphasizing future directions such as multi-table reasoning, long-context processing, and financial terminology comprehension. All our datasets, models [^1] https://huggingface.co/TheFinAI, and codes [^2] https://github.com/The-FinAI/Fino1 are publicly available. Furthermore, we introduce a leaderboard for benchmarking future datasets and models [^3] https://huggingface.co/spaces/TheFinAI/open-finllm-reasoning-leaderboard.
1 Introduction
Advancements in large language models (LLMs) have demonstrated remarkable performance across various natural language pocessing tasks, including content generationVan Veen et al. (2024); Zhang et al. (2024), language translation Xu et al. (2024), and sentiment analysis Xing (2024); Miah et al. (2024). More recently, reasoning-enhanced models such as OpenAI’s o1 and DeepSeek’s R1 have been developed to extend LLM capabilities in complex reasoning Guo et al. (2025); Jaech et al. (2024). These models exhibit significant improvements, particularly in mathematical and logical tasks that require complex reasoning Temsah et al. (2024); Zhong et al. (2024). However, despite their success in general reasoning, their performance in the financial domain remains largely unexplored.
Financial tasks inherently demand rigorous reasoning, requiring not only numerical computations but also a deep contextual understanding of financial terminology, regulations, and economic principles Xie et al. (2023, 2024a). Unlike general reasoning tasks, financial reasoning necessitates (1) comprehension of domain-specific terminology, (2) the ability to establish and manipulate relationships between financial concepts and numbers to derive accurate conclusions, and (3) an understanding of the structure and content of financial tables Xie et al. (2024b). Given the specialized nature of financial decision-making, assessing the reasoning capabilities of LLMs in this domain is crucial for bridging the gap between general AI reasoning and financial applications. Effective financial reasoning requires models to integrate textual, numerical, and structured data, highlighting the need for domain-specific adaptations to enhance AI-driven financial analysis.
In this study, we aim to comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks. Through this evaluation, we aim to provide insights into the strengths and limitations of existing reasoning models in financial applications, informing future advancements in financial AI research. We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks. (2) Impact across financial tasks – Analyzing how the reasoning enhancement process influences LLM performance across different types of financial tasks. (3) Scaling and performance gap – Investigating the differences between smaller reasoning models and large-scale LLMs in financial reasoning. (4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
We evaluated the performances of 16 large language models from the GPT Achiam et al. (2023), LLaMA Dubey et al. (2024), DeepSeek Liu et al. (2024), and Qwen Yang et al. (2024) families on three financial datasets, FinQA Chen et al. (2021), DM-Simplong Zhao et al. (2024), and XBRL-Math Wang et al. (2025), which incorporate financial text, tabular data, and equations as inputs. Our selection includes both general purpose and reasoning-enhanced models, such as DeepSeek-V3 Liu et al. (2024) and DeepSeek-R1 Guo et al. (2025), spanning from smaller models like LLaMA 3-8B-Instruct Dubey et al. (2024) to high-capacity models like GPT-o1 Jaech et al. (2024), ensuring a comprehensive assessment across different computational scales. These three financial reasoning datasets are designed to evaluate distinct aspects of financial understanding. Collectively, they assess models’ abilities in numerical reasoning, tabular data interpretation, financial terminology comprehension, long-context processing, and equation-based problem solving. Through this evaluation, our objective is to examine the capacity of LLMs to process financial concepts, reason over structured and unstructured financial data, and perform complex mathematical operations, offering deeper insight into their strengths and limitations in financial analysis tasks.
From experimental results and analysis, we have the following findings:
(1) General reasoning enhancement strategies, such as those employed in DeepSeek-R1 and GPT-o1, do not consistently improve performance in financial domain tasks. In fact, these models exhibit a performance decline compared to general-purpose models like DeepSeek-V3 and GPT-4o, suggesting challenges in understanding financial terminology and reasoning over long contexts and multi-table data.
(2) While improvements in pretraining methods and increased training data in general domains, as well as increasing model size all contributes to performance gains.For financial tasks, model performance tends to plateau at an upper boundary. Specifically, after reaching 70B or even 32B parameters, financial reasoning tasks show no significant improvement.
(3) The performance of different reasoning enhancement strategies in financial tasks can vary significantly. For instance, general-domain techniques like Chain-of-Thought and reflection-tuning in GPT-o1, designed for math and code, show no improvements in financial reasoning. In contrast, results show that DeepSeek and Qwen-Math incorporate enhance formula computations and structured data parsing using tool-integrated reasoning (TIR) and the process reward model (PRM), and knowledge distilled from general-purpose models. While these methods improve numerical accuracy in financial reports and market data, their impact on financial terminology comprehension, long-context modeling, and multi-table inference remains limited.
Based on the evaluation results, we developed a reasoning-enhanced language model specifically designed for financial tasks. We selected Llama-3.1-8B-Instruct Dubey et al. (2024) as the backbone model due to its superior performance among small-scale LLMs. To enhance its reasoning capabilities, we employed Chain-of-Thought (CoT) fine-tuning and reinforcement learning (RL), inspired by the four backtracking strategies from HuatuoGPT-o1 Chen et al. (2024) and the FinQA training dataset. To construct effective reasoning paths for fine-tuning, we leveraged GPT-4o, the best-performing closed-source model, as it achieved comparable results across all datasets. Experimental results indicate that even a simple fine-tuning approach using the reasoning paths from a single financial dataset, FinQA, led to significant performance improvements across multiple datasets. In some cases, our model’s performance approached that of advanced closed-source reasoning models, such as GPT-o3-mini and GPT-o1. These results demonstrate that reasoning enhancement tailored to financial domain tasks and datasets can significantly boost model performance in financial applications. In contrast, reasoning enhancement in general-domain tasks does not yield effective or consistent improvements for financial tasks, highlighting the necessity of domain-specific fine-tuning for optimal results.
To the best of our knowledge, we are the first to systematically explore the reasoning capabilities of enhanced LLMs in factual reasoning tasks across various financial inputs, including financial reports, tables, equations, and structured XBRL texts. Additionally, we investigate the effectiveness of reasoning enhancement strategies using financial data in adapting general-domain models for financial tasks. Overall, our results show that reasoning enhancement strategies designed for general domains do not consistently yield stable and effective improvements in financial reasoning tasks, whereas reasoning enhancement with financial data leads to significant performance gains. Furthermore, we highlight several key directions for advancing reasoning-enhanced LLMs in the financial domain, including enhancing models’ understanding of financial knowledge and terminology, improving their ability to handle multi-table reasoning and long financial contexts, and developing well-designed and diverse reasoning enhancement strategies to improve LLM performance in financial reasoning tasks.
2 Methods
2.1 Datasets
To evaluate the overall capabilities of existing reasoning-oriented and general LLMs in handling financial reasoning tasks, we select three tasks with different attributes:
(1) FINQA Chen et al. (2021): FinQA is a large-scale dataset designed to facilitate research in numerical reasoning within the financial domain. It consists of expert-annotated question-answer pairs and is distinctive in its focus on complex numerical reasoning, requiring the integration of both structured data (such as tables) and unstructured data (such as textual descriptions) found within financial reports. The dataset contains a vast number of financial terms, posing a challenge for models to understand domain-specific terminology and extract relevant information accurately.
(2) DocMath (simplong) Zhao et al. (2024): The DocMath-Eval dataset is a comprehensive benchmark designed to evaluate the numerical reasoning capabilities of LLMs, with a primary focus on long contexts consisit of specialized documents and tables. In particular, the simpllong subset is specifically designed to evaluate LLMs’ numerical reasoning abilities over long financial or specialized documents that contain multiple and multi-tiered tables. It challenges models to extract relevant information from extended contexts and perform reasoning across complex tabular structures.
(3) XBRL-Math Wang et al. (2025): This dataset is designed to evaluate the numerical reasoning capabilities of LLMs in the context of XBRL (eXtensible Business Reporting Language) filings. Different from regular financial reports, XBRL files are well structured that encode financial data using standardized taxonomies, ensuring consistency and interoperability across regulatory filings. It contains structured financial documents with US GAAP XBRL tags, equations, and multi-tiered numerical relationships. This dataset challenges models to extract, interpret, and reason across interconnected financial terms and formulas, such as APR = ((Fees + Interest) / Principal) × (365 / Days in Loan Term). Understanding XBRL taxonomy, hierarchical disclosures, and numerical dependencies is essential for accurate financial data analysis, making this dataset a benchmark for evaluating LLMs’ ability to handle complex financial reasoning tasks.
Dataset Size Data Types Average Token
FinQA 1100 Tables and Texts 1,128
DM-Simplong 100 Tables and Texts 4,330
XBRL-Math 90 Texts and Equations 397
Table 1: Overview of the datasets used in the study.
More details of the datasets are listed in Table 1. By conducting experiments on these three datasets, we aim to evaluate the model’s ability in various aspects, including understanding financial terms, extracting relevant numbers and entities from financial reports from different sources, and reasoning over long contexts and multiple tables.
2.2 Evaluated Models
To conduct a comprehensive evaluation of existing LLMs, we select models from the GPT, LLaMA, DeepSeek, and Qwen families, which have demonstrated effectiveness in various natural language processing tasks. Our selection includes models of different scales, ranging from small-sized models (8B) to large models (70B), as well as even larger ones like DeepSeek-R1 (approximately 700B). Additionally, we incorporate the most advanced closed-source models, such as GPT-4o and GPT-o1. Finally, all the models we use are listed following: (1) GPT-4o Hurst et al. (2024): OpenAI’s latest flagship model, featuring multimodal capabilities, improved efficiency, and real-time reasoning across text, image, and audio modalities. (2) GPT-o1 Jaech et al. (2024): A highly optimized variant within OpenAI’s GPT series, designed for enhanced reasoning and problem-solving with strong step-by-step thought processes. (3) GPT-o3-mini: A compact version of the o3 model, optimized for efficiency while maintaining robust language processing abilities. (4) DeepSeek-V3 Liu et al. (2024): DeepSeek’s Mixture-of-Experts (MoE) model with 671 billion parameters, activating 37 billion per token for efficient reasoning tasks. (5) DeepSeek-R1 Guo et al. (2025): DeepSeek’s first-generation reasoning model, achieving performance comparable to OpenAI’s o1 across math, code, and reasoning tasks. (6) Qwen2.5-72B-Instruct Yang et al. (2024): Alibaba’s open-source model demonstrating superior performance in Chinese and English tasks with advanced instruction tuning. (7) Qwen2.5-72B-Instruct-Math Yang et al. (2024): A specialized variant fine-tuned for mathematical reasoning, achieving strong results in complex problem-solving tasks. (8) Llama-series Models Dubey et al. (2024): This is the most widely adopted open-source large language model family, setting industry benchmarks for accessibility, efficiency, and performance. These models are trained on extensive datasets incorporating diverse multilingual and domain-specific data, leveraging high-quality pretraining and instruction tuning methodologies. In our experiments, we mainly focus on instruction finetune versions of these models ranging from 8B to 70B. (8) DeepSeek-R1-Distill Models Guo et al. (2025): This series of models is supervised fine-tuned on data distilled from DeepSeek-R1, leveraging its reasoning capabilities, with backbone models selected from Llama3.3-70B-Instruct, Qwen 2.5-32B, Qwen 2.5-14B, and Llama3.1-8B to balance efficiency and performance in instruction-following tasks.
Model Name Parameters Reasoning Enhanced Context Window Size Close/Open Source Reasoning Enhanced Training Data
GPT-4o Unknown No 128k Closed -
GPT-o1 Unknown Yes 128k Closed Public and properity data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini Unknown Yes 128k Closed Public and properity data
DeepSeek-V3 671B No 128k Open -
DeepSeek-R1 671B Yes 128k Open Cold-start data generation, post-processing data
Qwen2.5-72B-Instruct 72B No 128k Open -
Qwen2.5-72B-Instruct-Math 72B Yes 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
DeepSeek-R1-Distill-Llama-70B 70B Yes 128k Open Distilled from R1
Llama3-70B-Instruct 70B No 8k Open -
Llama3.1-70B-Instruct 70B No 128k Open -
Llama3.3-70B-Instruct 70B No 128k Open -
DeepSeek-R1-Distill-Qwen-32B 32B Yes 128k Open Distilled from R1
DeepSeek-R1-Distill-Qwen-14B 14B Yes 128k Open Distilled from R1
DeepSeek-R1-Distill-Llama-8B 8B Yes 128k Open Distilled from R1
Llama3-8B-Instruct 8B No 8k Open -
Llama3.1-8B-Instruct 8B No 128k Open -
Table 2: Summary of evaluated large language models, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies. MCTS refers to Monte Carlo Tree Search. CoT refers to chain-of-thought reasoning. TIR means tool-integrated reasoning.
2.3 Evaluation Settings
In each dataset, the LLMs are presented with a context consisting of financial reports, financial tables, equations, and mathematical symbols, along with a question. Insired by previous work Zhao et al. (2024), a generic prompt is used to instruct the model to answer the question based on the provided context. The models are required to extract relevant information and perform mathematical computations using numbers and concepts derived from the context. This necessitates strong logical reasoning abilities within financial contexts. The results are typically presented in a mathematical format, but due to variations in numerical representation—such as percentages and different rounding strategies—the final outputs can differ. Furthermore, reasoning-focused models, such as DeepSeek-R1 and GPT-o1, tend to provide a step-by-step reasoning process before arriving at the final answer. Inspired by DocMath, we employ an LLM-as-judge approach to evaluate model performance. First, an LLM (GPT-3.5-Turbo) is used to extract the final answer from the model’s output. Then, a mathematical comparison method is applied to determine whether the result is correct.
For large models like DeepSeek-V3 and DeepSeek-R1, or closed-source models like the GPT series, we conduct our experiments using APIs provided by togetherAI [^4] https://www.together.ai and OpenAI [^5] https://platform.openai.com. Specifically, we utilize GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06 for our evaluations. For other models with 72B parameters or fewer, we generate results using the VLLM framework on 4 A100 GPUs (80GB memory per GPU), with a maximum generation length of 512 tokens.
2.4 Fino1: LLMs for Financial Reasoning
To gain a deeper insight into how domain-specific reasoning enhancement benefits financial tasks compared to general reasoning enhancement, we developed Fino1, a reasoning-enhanced LLM specifically designed for financial tasks.
WTo enhance the domain-specific reasoning abilities of LLMs, we first construct high-quality training data with domain-specific reasoning paths based on the FinQA dataset. FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. This comprehensive dataset integrates both textual and tabular financial data, presenting a moderate level of difficulty, making it well-suited for training reasoning-based models.
For reasoning path generation, we utilize GPT-4o, as it demonstrates consistently strong performance across multiple datasets (as reported in Table 3), achieving results comparable to the highest-performing models on FinQA. Followed by HuatuoGPT-o1 Chen et al. (2024), we adopt a similar strategy to refine the CoT reasoning paths generated by GPT-4o. The refinement process involves the use of a verifier to assess the correctness of the generated CoT. If the verifier rejects the reasoning path, the model iteratively refines it by applying a strategy sampled from one of the following: (1) Backtracking – Revisiting previous steps to correct logical inconsistencies. (2) Exploring New Paths – Generating alternative reasoning routes when the current path fails. (3) Verification – Cross-checking intermediate steps to ensure logical soundness. (4) Correction – Making targeted adjustments to rectify errors. The CoT refinement process begins with an initial prompt to generate a reasoning path and an answer. The verifier then evaluates the response. If the answer is incorrect, one of the above strategies is applied iteratively to refine the CoT process until a correct and logically sound answer is generated.
2.4.1 Mode Training
After curating the reasoning data, we fine-tune Llama3.1-8B-Instruct, selected as the optimal backbone model based on the performance of 8B models. First, the curated dataset is used for supervised fine-tuning (SFT), where the question serves as input, and the reasoning path along with the final answer serves as output, enabling the model to learn structured financial reasoning. Next, we apply reinforcement learning with a reward model trained to assess the quality of generated reasoning paths. Inspired by prior work, we employ a verifier-based sparse reward mechanism and optimize the model using Proximal Policy Optimization (PPO). The verifier evaluates the correctness and coherence of the reasoning path; if an error is detected, the model iteratively refines the reasoning steps through backtracking, exploring alternative paths, verification, and correction. This two-stage approach enhances logical consistency, financial reasoning depth, and equation-solving accuracy, ultimately improving model performance across financial benchmarks.
During the SFT phase, the model was trained with a learning rate of 5e-6 for three epochs, using a maximum sequence length of 8192 tokens. LoRA (Low-Rank Adaptation) was applied to update parameters efficiently, reducing computational costs while maintaining performance. In the subsequent RL stage, the learning rate was lowered to 5e-7, and the model was trained for an additional three epochs. The reward model, previously trained in Chen et al. (2024), was employed to evaluate the quality of the model’s reasoning paths. This reward signal guided optimization through Proximal Policy Optimization (PPO), refining the model’s ability to generate accurate and coherent financial reasoning.
3 Results
3.1 Overall Results Analysis
The resutls of all these models on three dataset are listed in Table 3. Among all models, DeepSeek-R1 achieves the highest overall performance (68.93), driven by its strong XBRL-Math results, with its distilled variants, DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B, closely following. Among large-scale models, GPT-4o performs well across datasets but is outperformed by DeepSeek-R1 due to its lower XBRL-Math score. General-purpose 70B models, such as Qwen2.5-72B-Instruct and Llama3.3-70B-Instruct, achieve strong results, with Llama3.3-70B showing significant improvement over its predecessors. Surprisingly, these models also surpass the reasoning-focused GPT-o1 and approach GPT-o3-mini, suggesting that general-purpose reasoning strategies do not always translate effectively to financial tasks. In the 8B category, Fino1-8B (61.03) outperforms other models, demonstrating robust task optimization. Notably, reasoning-enhanced models exhibit mixed results: while they improve XBRL-Math scores, they tend to perform worse on FinQA and DM-Simplong. This suggests that different reasoning strategies yield varying benefits across financial reasoning tasks, emphasizing the importance of task-specific optimization.
Models FinQA DM-Simplong XBRL-Math Average
GPT-4o 72.49 60.00 72.22 68.24
GPT-o1 49.07 56.00 74.44 59.84
GPT-o3-mini 60.87 59.00 76.67 65.51
DeepSeek-V3 73.20 53.00 76.67 67.62
DeepSeek-R1 65.13 53.00 86.67 68.93
Qwen2.5-72B-Instruct 73.38 59.00 67.78 66.72
Qwen2.5-72B-Instruct-Math 69.74 42.00 83.33 65.69
DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 68.80
Llama3-70B-Instruct 58.92 41.00 56.67 52.20
Llama3.1-70B-Instruct 63.18 48.00 63.33 58.17
Llama3.3-70B-Instruct 68.15 54.00 70.00 64.05
DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 68.97
DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 63.90
DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 53.36
Llama3-8B-Instruct 41.97 29.00 48.89 39.95
Llama3.1-8B-Instruct 54.13 34.00 62.22 50.12
Fino1-8B 60.87 40.00 82.22 61.03
Table 3: Performance of different LLMs on three tested financial datasets.
From our results, we identify several key insights regarding the performance of large language models (LLMs) in financial reasoning tasks:
(1) General reasoning enhancements do not necessarily translate to improvements in financial tasks. While models fine-tuned for general reasoning, such as GPT-o1 and DeepSeek-R1, demonstrate strong performance in mathematical and logical reasoning tasks, they do not consistently outperform other models in financial benchmarks. For example, GPT-o1 scores 49.07 in FinQA and 56.0 in DM-Simplong, trailing behind models like GPT-4o that are not explicitly optimized for general reasoning. Also, DeepSeek-R1 also under performan V3 in these tasks. This suggests that financial reasoning requires more than just enhanced logical and mathematical capabilities; it demands a deep understanding of domain-specific terminology, structured financial statements, and numerical patterns. One possible explanation for this phenomenon is that reasoning-focused fine-tuning, which emphasizes general problem-solving strategies, may inadvertently weaken a model’s retention of financial knowledge. This is particularly relevant for tasks requiring precise comprehension of financial terms, balance sheet structures, and industry-specific metrics, where models without targeted financial fine-tuning may struggle. Additionally, reasoning-enhanced models might prioritize general logical inference over domain-specific factual recall, which could lead to a decrease in performance when interpreting specialized financial contexts. These findings indicate that while general reasoning improvements enhance broad problem-solving skills, they do not necessarily translate into domain-specific advantages, highlighting the need for specialized financial training in LLMs.
(2) Different reasoning enhancement strategies yield varying improvements depending on the financial reasoning task format. By comparing the performance of GPT-4o vs. GPT-o1, Qwen2.5-72B-Instruct vs. Qwen2.5-72B-Instruct-Math, and Llama models vs. the DeepSeek-R1 distilled Llama models, we observe that reasoning-enhanced models do not always improve financial task performance. GPT-o1 underperforms GPT-4o in FinQA (49.07 vs. 72.49) and DM-Simplong (56.00 vs. 60.00), suggesting that its reasoning strategies, such as chain of thought and reflection-tuning, are optimized for general problem-solving rather than financial-specific logic. Similarly, Qwen2.5-72B-Instruct-Math, despite achieving a substantial XBRL-Math boost (67.78 → 83.33), experiences sharp declines in FinQA (69.74 → 42.00) and DM-Simplong (59.00 → 42.00), indicating that mathematical reasoning enhancements help structured financial data but do not generalize well to broader financial tasks. In the 8B category, DeepSeek-R1-Distill-Llama-8B (53.36) outperforms Llama3.1-8B-Instruct (50.12) and Llama3-8B-Instruct (39.95), excelling in XBRL-Math (81.11) but showing limited improvements in other financial tasks.
These results suggest that different reasoning strategies target distinct aspects of financial reasoning. DeepSeek-R1 and Qwen2.5-72B-Instruct-Math demonstrate strong performance in symbolic and mathematical reasoning, excelling in structured equation-based tasks. However, they do not generalize well to financial question-answering tasks like FinQA and DM-Simplong. Similarly, reasoning augmentations in GPT-o1 fail to provide consistent improvements, likely due to the absence of relevant financial reasoning data, highlighting the challenge of adapting broad reasoning techniques to financial logic.
(3) Scaling model size does not always lead to performance gains in financial tasks. While larger models typically achieve better results in standard NLP tasks, their performance in financial tasks does not always scale proportionally with size. For instance, DeepSeek-70B (66.73 in FinQA, 53 in DM-Simplong, and 86.67 in XBRL-Math) performs similarly to its smaller distilled variant DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite the latter having less than half the parameters. A broader comparison, including Qwen2.5-72B-Instruct, DeepSeek-R1-Distill-Llama-70B, GPT-4o, DeepSeek-V3, and DeepSeek-R1, reveals that performance stabilizes once model size reaches approximately 70B parameters. One possible explanation for this trend is the nature of XBRL-Math, which consists of structured financial tables and numerical equations that prioritize precise mathematical computations over complex language-based reasoning. In such structured, numerically driven tasks, smaller models can achieve competitive performance by leveraging arithmetic operations and pattern recognition rather than advanced language modeling. This suggests that the advantage of larger models diminishes in financial decision-making tasks where structured data comprehension and numerical reasoning play a more crucial role than language-based inference.
(4) Improving Pre-training Data and Strategies in General Domains Also Benefits Financial Tasks. Compared with Llama 3-8B-Instruct, Llama 3.1-8B-Instruct demonstrates a significant performance boost, increasing its average score from 39.95 to 50.12, with substantial improvements in XBRL-Math (62.22 vs. 48.89) and FinQA (54.13 vs. 41.97). Similarly, among the 70B-Instruct models, Llama 3.1-70B-Instruct outperforms Llama 3-70B-Instruct (58.17 vs. 52.20), with the largest improvement in DM-Simplong (48.00 vs. 41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05 vs. 58.17), excelling in FinQA (68.15 vs. 63.18) and XBRL-Math (70.00 vs. 63.33), highlighting the benefits of dataset expansion and refined training strategies.
More specifically, Llama 3.1 incorporates a more extensive dataset, covering over 15 trillion tokens across multiple languages and domains, leading to notable performance gains. Llama 3.3 further benefits from advancements in post-training techniques, such as online preference optimization, which enhances core performance while significantly reducing training and inference costs. These refined optimization strategies in pre-training not only improve general capabilities but also enhance performance in domain-specific tasks.
(5) Existing reasoning models struggle with long financial contexts and cross-table correlations. Many financial reasoning tasks require models to extract and synthesize information from lengthy reports, structured tables, and numerical equations. However, performance discrepancies across datasets suggest that current LLMs struggle with this challenge. For example, models that perform well in FinQA, which primarily involves textual financial reasoning, do not necessarily excel in DM-Simplong, which integrates numerical and symbolic information. GPT-4o achieves 72.49 in FinQA but drops to 60.0 in DM-Simplong, indicating difficulties in integrating mathematical symbols with textual reasoning. Furthermore, models that score highly in XBRL-Math, such as DeepSeek-R1-Distill-Qwen-32B (84.44) and DeepSeek-R1 (86.67), highlight a growing but incomplete capability to process structured financial tables. Notably, all reasoning-enhanced models fail to show improvements in handling long financial contexts and multi-table correlations. This is particularly concerning given that financial reports are inherently long and include complex interlinked tables. Though smaller models perform well in structured inputs like XBRL-Math, their performance drops significantly—by nearly 30%—when handling long contexts. For example, DeepSeek-R1-Distill-Llama-8B achieves 81.11 in XBRL-Math but drops to 33.0 in DM-Simplong. This suggests that while smaller models can recognize structured patterns, they struggle with long-range dependencies and multi-step reasoning. Addressing these limitations requires improvements in long-context retention, structured data comprehension, and multi-modal document understanding to enhance financial reasoning capabilities.
(6) Reasoning enhancement basd on financial data improve performance of all datasets. Compared with Llama 3.1-8B-Instruct, DeepSeek-R1-Distill-Llama-8B, which incorporates reasoning enhancement data distilled from DeepSeek-R1, shows only a modest improvement (50.12 to 53.36). However, our model, Fino1-8B, trained with reasoning paths distilled from GPT-4o using a backtracking strategy to iteratively refine the reasoning process, achieves a substantial performance boost across all datasets, with an average score of 61.03—a 10.91-point increase over Llama 3.1-8B-Instruct. Notably, Fino1-8B outperforms its counterparts in FinQA (60.87 vs. 54.13), DM-Simplong (40.00 vs. 34.00), and particularly in XBRL-Math (82.22 vs. 62.22), demonstrating the effectiveness of reasoning-driven training.
It is important to note that our training utilized reasoning paths generated exclusively from FinQA, yet the model exhibited improvements across all datasets. This suggests that financial domain-specific reasoning enhancement data can significantly improve models’ comprehensive understanding of financial terminology, long-context comprehension, table-based reasoning, and equation interpretation. Furthermore, these results highlight the effectiveness of a combination of reasoning path-building strategies, including backtracking, exploring new paths, verification, and correction, in financial tasks. By refining reasoning paths through iterative corrections, these strategies enhance logical consistency, computational accuracy, and overall robustness, demonstrating a promising approach for advancing financial reasoning capabilities in language models.
Overall, these insights demonstrate that financial reasoning presents unique challenges distinct from general NLP and mathematical reasoning. The variability in performance across different tasks underscores the necessity for more financially specialized training methodologies and architectural improvements to enhance financial comprehension, structured data processing, and multi-step numerical reasoning.
3.2 Error Analysis
To further investigate why reasoning-enhanced models like DeepSeek-R1 and GPT-o1 underperform in financial tasks compared to general-purpose models such as GPT-4o and DeepSeek-V3, we conduct an error analysis using the FinQA dataset since these reasoning-enhanced models exhibit a significant performance drop on FinQA.
We focus on cases where DeepSeek-R1 produces incorrect answers while DeepSeek-V3 generates correct ones. Two common error patterns emerge. First, DeepSeek-R1 tends to over-reason and does not strictly adhere to instructions, as illustrated in Figure 1. In this case, the question requires a sum, but the model provides excessive details by breaking down the calculation instead of directly summing the values as instructed.
Second, compared to DeepSeek-V3, DeepSeek-R1 appears to lack financial sensitivity. For instance, in Figure 2, it fails to recognize critical financial nuances, leading to an incorrect response. Specifically, DeepSeek-R1 misinterprets the concept of "average share price" in the UK Employee Share Purchase Plan by incorrectly identifying it as the employee purchase price (\$4,078 per share) rather than the actual compensation expense per share (\$719 per share). This mistake stems from an inability to correctly differentiate between compensation expense per share and the full fair market value (FMV) calculation, resulting in an overestimated figure. In contrast, DeepSeek-V3 correctly understands that the compensation expense per share represents the direct financial impact per issued share, allowing it to provide the accurate answer. This highlights a fundamental gap in DeepSeek-R1’s financial reasoning capabilities, particularly in recognizing the accounting conventions used to report share-based compensation.
Figure 1: Error case 1.
Figure 2: Error case 2.
3.3 Future Direction
Based on the above findings, we observe that developing a powerful reasoning model for financial tasks presents unique challenges distinct from general-domain reasoning. Our analysis suggests three key areas for improvement in financial reasoning models:
(1) Enhancing models’ understanding of financial knowledge and terminology. While reasoning-enhanced models demonstrate strong performance in general domains, they often struggle with financial tasks due to a lack of domain-specific knowledge. Future work should explore strategies that incorporate financial corpora, structured financial reports, and regulatory documents to improve models’ comprehension of domain-specific terminology and contextual nuances.
(2) Improving models’ ability to handle multi-table reasoning and long financial contexts. Our results highlight that even large models exhibit significant performance drops when processing long financial reports and cross-referencing multiple tables. To address this, future reasoning stratigies should be focused on inferring logics among tables or menorization mechanisms to enhance long-context understanding.
(3) Deeper analysis of reasoning-enhanced strategies in models like GPT-o1 and DeepSeek-R1. Although GPT-o1 and DeepSeek-R1 are designed with advanced reasoning enhancements, they fail to show consistent improvements across financial tasks. GPT-o1 utilizes "chain-of-thought" prompting, generating intermediate reasoning steps to enhance structured problem-solving. However, this reasoning strategy does not lead to improvements in any of the financial benchmarks, as GPT-o1 underperforms compared to GPT-4o across all tasks, scoring 49.07 in FinQA and 56.0 in DM-Simplong. In contrast, DeepSeek-R1 introduces self-reflection capabilities through reinforcement learning, enabling the model to iteratively evaluate and refine its reasoning processes without supervised data. While this approach helps DeepSeek-R1 achieve strong performance in XBRL-Math (86.67), indicating improved structured numerical reasoning, it also leads to a drop in FinQA (65.13), suggesting challenges in financial text-based reasoning. These results highlight a fundamental difference between the two models — GPT-o1’s structured reasoning approach does not translate to domain-specific improvements, whereas DeepSeek-R1’s reinforcement learning strategy enhances structured data processing but remains limited in free form textual financial reasoning. Understanding these trade-offs is essential for designing reasoning strategies tailored to financial tasks.
Developing robust financial reasoning models requires both domain adaptation and architectural advancements to bridge the gap between general reasoning capabilities and financial-specific challenges. Future work should focus on integrating financially specialized pretraining, multi-table comprehension techniques, and long-context optimization to build models capable of handling complex financial documents and reasoning tasks effectively.
4 Related Work
Our work builds upon three main research directions: the development of general-purpose and reasoning-enhanced language models, applications of LLMs in finance, and studies on reasoning capabilities in financial tasks.
4.1 General-Purpose and Reasoning-Enhanced LLMs
Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning. Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting Wei et al. (2022), self-reflection Zhang et al. (2023), and reinforcement learning from human feedback Ouyang et al. (2022). These models have shown superior performance in mathematical reasoning Lewkowycz et al. (2022), symbolic manipulation Drori et al. (2023), and logical inference tasks Chen et al. (2023).
4.2 LLMs in Financial Applications
The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis Yang et al. (2020), market sentiment prediction Yang et al. (2023), and automated trading strategies Kou et al. (2024). Recent work has focused on developing finance-specific models like BloombergGPT Wu et al. (2023), FinGPT Yang et al. (2023) and PIXIU Xie et al. (2023), OpenFinLLMs Xie et al. (2024b), which are pretrained on financial corpora to better understand domain-specific terminology and concepts. These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
4.3 Reasoning Capabilities in Financial Tasks
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al. (2021), mathematical problem-solving in financial contexts Srivastava et al. (2024), and structured data interpretation Liu et al. (2025). These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
5 Conclusion
In this study, we evaluated the performance of 16 recent reasoning-augmented LLMs and general-purpose LLMs on financial tasks, analyzing their capabilities in text understanding, numerical reasoning, and structured data processing. Our results indicate that while reasoning-enhanced models demonstrate strong performance in general tasks, they often fail to generalize effectively to financial reasoning due to insufficient domain adaptation. Additionally, models optimized for structured mathematical reasoning perform well in numerical tasks like XBRL-Math but struggle with financial text comprehension and long-context processing, highlighting the need for financial domain-specific adaptations. To address these limitations, we presented Fino1, a reasoning-enhanced LLMs based on Llama-3.1-8B-Instruct, leveraging reasoning paths curated from GPT-4o using only FinQA, a dataset with 5,000 financial reasoning samples. Our results demonstrate that, despite being trained on a single dataset, Fino1 achieved an overall performance improvement of approximately 10% across three financial benchmarks. This significant improvement underscores the potential of financial domain-specific data in enhancing LLMs’ financial reasoning capabilities. Future work can focus on enhancing financial knowledge adaptation, improving multi-table reasoning, and handling long-context financial documents. Additionally, refining domain-specific reasoning strategies through structured data reasoning, retrieval-augmented approaches, and domain-aware pretraining will be critical for bridging the gap between general reasoning capabilities and financial-specific challenges.
Limitations
Our study has several limitations that highlight areas for future improvement. First, our model scale is limited, as we only fine-tuned an 8B model (Fino1), while larger models (e.g., 70B) could potentially benefit more from reasoning enhancements. Second, our evaluation scope is restricted, covering only three financial reasoning tasks (FinQA, DM-Simplong, and XBRL-Math), which do not fully capture the breadth of financial NLP applications such as forecasting, financial sentiment analysis, and fraud detection. Third, our fine-tuning relies on a single dataset, FinQA, for reasoning path construction, limiting the model’s exposure to different financial reasoning patterns; incorporating additional datasets could improve generalization. Fourth, our reasoning path construction approach is simplified, as we generate paths using a single method (GPT-4o), whereas exploring multiple reasoning path generation strategies—such as ensemble approaches or human-annotated paths—could lead to more robust financial reasoning capabilities.
Ethics Statement
While our study demonstrates significant advancements in financial reasoning with LLMs, Fino1 remains a research prototype and is not yet suitable for real-world financial applications. Our model inherits the well-documented limitations of large language models, including hallucinations, sensitivity to input phrasing, and potential biases in financial data sources, which could affect reliability in critical financial contexts. Furthermore, Fino1 has not undergone rigorous testing for high-stakes financial decision-making, such as investment strategies, regulatory compliance, or risk assessment, where incorrect outputs could lead to severe consequences. Additionally, the model’s reliance on a single dataset (FinQA) and a single reasoning path construction method (GPT-4o) limits its robustness and adaptability to the complex and evolving nature of financial environments. As with all AI-driven financial models, careful human oversight remains essential, and future work must focus on enhancing factual consistency, mitigating biases, improving domain adaptation, and ensuring alignment with regulatory frameworks before such models can be considered for real-world financial deployment.

### v2

@@
 Abstract
-Recent advancements in large language models (LLMs) have shown strong general reasoning abilities, yet their effectiveness in financial reasoning remains underexplored.
-In this study, we comprehensively evaluate 16 powerful reasoning and general LLMs on three complex financial tasks involving financial text, tabular data, and equations, assessing numerical reasoning, tabular interpretation, financial terminology comprehension, long-context processing, and equation-based problem solving.
-Our results show that while better datasets and pretraining improve financial reasoning, general enhancements like CoT fine-tuning do not always yield consistent gains.
-Moreover, all reasoning strategies face challenges in improving performance on long-context and multi-table tasks.
-To address these limitations, we develop a financial reasoning-enhanced model based on Llama-3.1-8B-Instruct, by CoT fine-tuning and reinforcement learning with domain-specific reasoning paths.
-Even with simple fine-tuning with one financial dataset, our model achieves a consistent 10% performance improvement across tasks, surpassing all 8B models and even Llama3-70B-Instruct and Llama3.1-70B-Instruct on average.
-Our results highlight the need for domain-specific adaptations in financial tasks, emphasizing future directions such as multi-table reasoning, long-context processing, and financial terminology comprehension.
-All our datasets, models [^1] https://huggingface.co/TheFinAI, and codes [^2] https://github.com/The-FinAI/Fino1 are publicly available.
-Furthermore, we introduce a leaderboard for benchmarking future datasets and models [^3] https://huggingface.co/spaces/TheFinAI/open-finllm-reasoning-leaderboard.
+While large language models (LLMs) have shown strong general reasoning capabilities, their effectiveness in financial reasoning—which is crucial for real-world financial applications—remains underexplored.
+In this study, we conduct a comprehensive evaluation of 24 state-of-the-art general and reasoning-focused LLMs across four complex financial reasoning tasks involving financial text, tabular data, and equations.
+We assess key capabilities such as numerical reasoning, tabular interpretation, financial terminology comprehension, long-context understanding, and equation-based problem solving.
+Our analysis reveals that while data quality and pretraining contribute to performance, general techniques like chain-of-thought (CoT) fine-tuning offer limited gains in financial tasks.
+To address this, we propose two domain-adapted models, Fino1-8B and Fino1-14B, trained with CoT fine-tuning and reinforcement learning using domain-specific reasoning paths.
+Our models are trained on a carefully curated dataset integrating high-quality examples from diverse sources, covering financial reports, tables, equations, and structured XBRL texts.
+Despite limited training data, they achieve an 7–9% performance improvement, outperforming several advanced LLMs—including GPT-o1, GPT-o3-mini, GPT-4.5, and comparable with DeepSeek models (V3 and R1)—demonstrating strong practical value in resource-constrained scenarios.
+Our findings highlight the need for domain-specific adaptations in financial reasoning, and we release all datasets, models, and code for future research [^1] https://github.com/The-FinAI/Fino1.
 1 Introduction
@@
 However, despite their success in general reasoning, their performance in the financial domain remains largely unexplored.
-Financial tasks inherently demand rigorous reasoning, requiring not only numerical computations but also a deep contextual understanding of financial terminology, regulations, and economic principles Xie et al.
+Financial reasoning is essential in the financial domain, where professionals must extract key insights from lengthy reports and multi-table documents.
+It requires not only numerical computation but also a deep understanding of financial terminology, regulations, and economic concepts Xie et al.
 (2023, 2024a).
-Unlike general reasoning tasks, financial reasoning necessitates (1) comprehension of domain-specific terminology, (2) the ability to establish and manipulate relationships between financial concepts and numbers to derive accurate conclusions, and (3) an understanding of the structure and content of financial tables Xie et al.
+Unlike general reasoning tasks, financial reasoning involves (1) domain-specific language comprehension, (2) reasoning over financial concepts and numbers, and (3) interpreting structured data such as financial tables Xie et al.
 (2024b).
-Given the specialized nature of financial decision-making, assessing the reasoning capabilities of LLMs in this domain is crucial for bridging the gap between general AI reasoning and financial applications.
-Effective financial reasoning requires models to integrate textual, numerical, and structured data, highlighting the need for domain-specific adaptations to enhance AI-driven financial analysis.
-In this study, we aim to comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks.
-Through this evaluation, we aim to provide insights into the strengths and limitations of existing reasoning models in financial applications, informing future advancements in financial AI research.
+Evaluating LLMs in this context is critical for closing the gap between general AI reasoning and real-world financial applications.
+To address this gap, we firstly comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks, to provide insights into the strengths and limitations of existing reasoning models in financial applications.
 We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks.
@@
 (4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
-We evaluated the performances of 16 large language models from the GPT Achiam et al.
+We evaluated 24 LLMs from the GPT Achiam et al.
 (2023), LLaMA Dubey et al.
@@
 (2024), and Qwen Yang et al.
-(2024) families on three financial datasets, FinQA Chen et al.
+(2024) families on four financial reasoning benchmarks: FinQA Chen et al.
 (2021), DM-Simplong Zhao et al.
-(2024), and XBRL-Math Wang et al.
-(2025), which incorporate financial text, tabular data, and equations as inputs.
-Our selection includes both general purpose and reasoning-enhanced models, such as DeepSeek-V3 Liu et al.
-(2024) and DeepSeek-R1 Guo et al.
-(2025), spanning from smaller models like LLaMA 3-8B-Instruct Dubey et al.
-(2024) to high-capacity models like GPT-o1 Jaech et al.
-(2024), ensuring a comprehensive assessment across different computational scales.
-These three financial reasoning datasets are designed to evaluate distinct aspects of financial understanding.
-Collectively, they assess models’ abilities in numerical reasoning, tabular data interpretation, financial terminology comprehension, long-context processing, and equation-based problem solving.
-Through this evaluation, our objective is to examine the capacity of LLMs to process financial concepts, reason over structured and unstructured financial data, and perform complex mathematical operations, offering deeper insight into their strengths and limitations in financial analysis tasks.
-From experimental results and analysis, we have the following findings:
-(1) General reasoning enhancement strategies, such as those employed in DeepSeek-R1 and GPT-o1, do not consistently improve performance in financial domain tasks.
-In fact, these models exhibit a performance decline compared to general-purpose models like DeepSeek-V3 and GPT-4o, suggesting challenges in understanding financial terminology and reasoning over long contexts and multi-table data.
-(2) While improvements in pretraining methods and increased training data in general domains, as well as increasing model size all contributes to performance gains.For financial tasks, model performance tends to plateau at an upper boundary.
-Specifically, after reaching 70B or even 32B parameters, financial reasoning tasks show no significant improvement.
-(3) The performance of different reasoning enhancement strategies in financial tasks can vary significantly.
-For instance, general-domain techniques like Chain-of-Thought and reflection-tuning in GPT-o1, designed for math and code, show no improvements in financial reasoning.
-In contrast, results show that DeepSeek and Qwen-Math incorporate enhance formula computations and structured data parsing using tool-integrated reasoning (TIR) and the process reward model (PRM), and knowledge distilled from general-purpose models.
-While these methods improve numerical accuracy in financial reports and market data, their impact on financial terminology comprehension, long-context modeling, and multi-table inference remains limited.
-Based on the evaluation results, we developed a reasoning-enhanced language model specifically designed for financial tasks.
-We selected Llama-3.1-8B-Instruct Dubey et al.
-(2024) as the backbone model due to its superior performance among small-scale LLMs.
-To enhance its reasoning capabilities, we employed Chain-of-Thought (CoT) fine-tuning and reinforcement learning (RL), inspired by the four backtracking strategies from HuatuoGPT-o1 Chen et al.
-(2024) and the FinQA training dataset.
-To construct effective reasoning paths for fine-tuning, we leveraged GPT-4o, the best-performing closed-source model, as it achieved comparable results across all datasets.
-Experimental results indicate that even a simple fine-tuning approach using the reasoning paths from a single financial dataset, FinQA, led to significant performance improvements across multiple datasets.
-In some cases, our model’s performance approached that of advanced closed-source reasoning models, such as GPT-o3-mini and GPT-o1.
-These results demonstrate that reasoning enhancement tailored to financial domain tasks and datasets can significantly boost model performance in financial applications.
-In contrast, reasoning enhancement in general-domain tasks does not yield effective or consistent improvements for financial tasks, highlighting the necessity of domain-specific fine-tuning for optimal results.
-To the best of our knowledge, we are the first to systematically explore the reasoning capabilities of enhanced LLMs in factual reasoning tasks across various financial inputs, including financial reports, tables, equations, and structured XBRL texts.
-Additionally, we investigate the effectiveness of reasoning enhancement strategies using financial data in adapting general-domain models for financial tasks.
-Overall, our results show that reasoning enhancement strategies designed for general domains do not consistently yield stable and effective improvements in financial reasoning tasks, whereas reasoning enhancement with financial data leads to significant performance gains.
-Furthermore, we highlight several key directions for advancing reasoning-enhanced LLMs in the financial domain, including enhancing models’ understanding of financial knowledge and terminology, improving their ability to handle multi-table reasoning and long financial contexts, and developing well-designed and diverse reasoning enhancement strategies to improve LLM performance in financial reasoning tasks.
+(2024a), XBRL-Math Wang et al.
+(2025) and DM-Complong.
+These datasets span financial texts, tables, and equations, enabling evaluation of models in numerical reasoning, domain-specific terminology comprehension, long-context processing, and equation-based reasoning tasks.
+Our evaluation encompasses both general-purpose and reasoning-enhanced models, ranging from smaller-scale models (e.g., LLaMA 3-8B-InstructDubey et al.
+(2024)) to high-capacity models (e.g., GPT-o1Jaech et al.
+(2024)), as well as reasoning-augmented models incorporating strategies such as data sampling and financial Chain-of-Thought (CoT) construction.
+Experimental results reveal three key insights: (1) Reasoning-enhanced models like DeepSeek-R1 and GPT-o1 do not consistently outperform general-purpose models on financial tasks, likely due to challenges related to domain-specific terminology and multi-table reasoning; (2) Performance gains from larger model sizes and improved pretraining tend to plateau beyond 32B or 70B parameters;
+To develop reasoning models tailored for financial domain, we build novel financial reasoning-enhanced large language models Fino1-8B and Fino1-14B based on LLaMA-3.1-8B-InstructDubey et al.
+(2024) and Qwen2.5-14B-InstructYang et al.
+(2024), chosen for their strong performance among smaller-scale LLMs.
+To enhance reasoning capabilities, we apply CoT fine-tuning and reinforcement learning (RL), inspired by the training strategies of HuatuoGPT-o1 Chen et al.
+(2024).
+We leverage a diverse set of high quality financial reasoning datasets—including FinQA Chen et al.
+(2022), TATQA Zhu et al.
+(2021), DocMath (Simpshort and Compshort) Zhao et al.
+(2024b) , DocFinQA Reddy et al.
+(2024a), and BizBench-QA Krumdick et al.
+(2024)—covering tasks such as long-context question answering, complex reasoning over financial documents, and table-text integration.
+We construct high-quality reasoning paths using GPT-4o, combined with a novel data filtering process that selects more challenging samples for training—particularly for the 14B model—ensuring the model learns from difficult, domain-specific reasoning cases rather than generic QA pairs.
+Experimental results show that this targeted training strategy enables our models to surpass the performance of advanced closed-source models such as GPT-o3-mini and GPT-o1, and approach most powerful LLMs like DeepSeek-V3, DeepSeek-R1, and GPT-4o, demonstrating that domain-specific reasoning-enhanced training with limited, high-quality data can be highly effective.
+To the best of our knowledge, this work presents the first comprehensive study of reasoning-enhanced LLMs for financial tasks involving diverse inputs such as financial documents, tables, equations, and structured XBRL texts.
+First, we conduct a systematic evaluation of 24 general-purpose and reasoning-enhanced LLMs across multiple financial reasoning benchmarks.
+Second, we introduce Fino1, the first suite of domain-specific, reasoning-enhanced LLMs, trained using GPT-4o-distilled reasoning paths with a novel data curation and filtering strategy.
+Third, our results show that domain-specific reasoning training with limited high-quality data consistently outperforms general-domain strategies and even surpasses larger closed-source models.
+We release all models, datasets, and code to support future research in financial AI.
 2 Methods
 2.1 Datasets
-To evaluate the overall capabilities of existing reasoning-oriented and general LLMs in handling financial reasoning tasks, we select three tasks with different attributes:
+To evaluate the financial reasoning capabilities of reasoning-oriented and general LLMs, we select four diverse benchmark tasks.
 (1) FINQA Chen et al.
-(2021): FinQA is a large-scale dataset designed to facilitate research in numerical reasoning within the financial domain.
-It consists of expert-annotated question-answer pairs and is distinctive in its focus on complex numerical reasoning, requiring the integration of both structured data (such as tables) and unstructured data (such as textual descriptions) found within financial reports.
-The dataset contains a vast number of financial terms, posing a challenge for models to understand domain-specific terminology and extract relevant information accurately.
-(2) DocMath (simplong) Zhao et al.
-(2024): The DocMath-Eval dataset is a comprehensive benchmark designed to evaluate the numerical reasoning capabilities of LLMs, with a primary focus on long contexts consisit of specialized documents and tables.
-In particular, the simpllong subset is specifically designed to evaluate LLMs’ numerical reasoning abilities over long financial or specialized documents that contain multiple and multi-tiered tables.
-It challenges models to extract relevant information from extended contexts and perform reasoning across complex tabular structures.
+(2021) is a large-scale dataset that emphasizes complex numerical reasoning in financial reports, requiring the integration of structured tables and unstructured text, along with a deep understanding of domain-specific terminology.
+(2) DocMath (Simplong, Complong) Zhao et al.
+(2024a) evaluates LLMs’ numerical reasoning abilities over long, specialized documents with multi-tiered tables, testing their capacity to extract and reason over extended contexts.
+The simplong subset contains relatively easier questions, while the complong subset includes more complex questions that require advanced reasoning across tables and context.
 (3) XBRL-Math Wang et al.
-(2025): This dataset is designed to evaluate the numerical reasoning capabilities of LLMs in the context of XBRL (eXtensible Business Reporting Language) filings.
-Different from regular financial reports, XBRL files are well structured that encode financial data using standardized taxonomies, ensuring consistency and interoperability across regulatory filings.
-It contains structured financial documents with US GAAP XBRL tags, equations, and multi-tiered numerical relationships.
-This dataset challenges models to extract, interpret, and reason across interconnected financial terms and formulas, such as APR = ((Fees + Interest) / Principal) × (365 / Days in Loan Term).
-Understanding XBRL taxonomy, hierarchical disclosures, and numerical dependencies is essential for accurate financial data analysis, making this dataset a benchmark for evaluating LLMs’ ability to handle complex financial reasoning tasks.
+(2025) targets reasoning over structured financial data in XBRL filings, which involve standardized taxonomies, hierarchical disclosures, and formulas, challenging models to interpret intricate numerical relationships.
 Dataset Size Data Types Average Token
@@
 DM-Simplong 100 Tables and Texts 4,330
+DM-Complong 300 Tables and Texts 39,983
 XBRL-Math 90 Texts and Equations 397
@@
 2.2 Evaluated Models
-To conduct a comprehensive evaluation of existing LLMs, we select models from the GPT, LLaMA, DeepSeek, and Qwen families, which have demonstrated effectiveness in various natural language processing tasks.
-Our selection includes models of different scales, ranging from small-sized models (8B) to large models (70B), as well as even larger ones like DeepSeek-R1 (approximately 700B).
-Additionally, we incorporate the most advanced closed-source models, such as GPT-4o and GPT-o1.
-Finally, all the models we use are listed following: (1) GPT-4o Hurst et al.
-(2024): OpenAI’s latest flagship model, featuring multimodal capabilities, improved efficiency, and real-time reasoning across text, image, and audio modalities.
-(2) GPT-o1 Jaech et al.
-(2024): A highly optimized variant within OpenAI’s GPT series, designed for enhanced reasoning and problem-solving with strong step-by-step thought processes.
-(3) GPT-o3-mini: A compact version of the o3 model, optimized for efficiency while maintaining robust language processing abilities.
-(4) DeepSeek-V3 Liu et al.
-(2024): DeepSeek’s Mixture-of-Experts (MoE) model with 671 billion parameters, activating 37 billion per token for efficient reasoning tasks.
-(5) DeepSeek-R1 Guo et al.
-(2025): DeepSeek’s first-generation reasoning model, achieving performance comparable to OpenAI’s o1 across math, code, and reasoning tasks.
-(6) Qwen2.5-72B-Instruct Yang et al.
-(2024): Alibaba’s open-source model demonstrating superior performance in Chinese and English tasks with advanced instruction tuning.
-(7) Qwen2.5-72B-Instruct-Math Yang et al.
-(2024): A specialized variant fine-tuned for mathematical reasoning, achieving strong results in complex problem-solving tasks.
-(8) Llama-series Models Dubey et al.
-(2024): This is the most widely adopted open-source large language model family, setting industry benchmarks for accessibility, efficiency, and performance.
-These models are trained on extensive datasets incorporating diverse multilingual and domain-specific data, leveraging high-quality pretraining and instruction tuning methodologies.
-In our experiments, we mainly focus on instruction finetune versions of these models ranging from 8B to 70B.
-(8) DeepSeek-R1-Distill Models Guo et al.
-(2025): This series of models is supervised fine-tuned on data distilled from DeepSeek-R1, leveraging its reasoning capabilities, with backbone models selected from Llama3.3-70B-Instruct, Qwen 2.5-32B, Qwen 2.5-14B, and Llama3.1-8B to balance efficiency and performance in instruction-following tasks.
-Model Name Parameters Reasoning Enhanced Context Window Size Close/Open Source Reasoning Enhanced Training Data
+[…]
+The evaluated models are summarized as follows: GPT-4o Hurst et al.
+(2024): OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning.
+GPT-o1 Jaech et al.
+(2024): A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability.
+GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding.
+DeepSeek-V3 Liu et al.
+(2024): A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning.
+DeepSeek-R1 Guo et al.
+(2025): A first-generation reasoning model with competitive performance on math, code, and reasoning tasks.
+Qwen2.5-72B-Instruct Yang et al.
+(2024): A strong multilingual model with advanced instruction tuning.
+Qwen2.5-72B-Instruct-Math: A math-specialized variant fine-tuned for complex problem solving.
+LLaMA-series Dubey et al.
+(2024): Widely adopted open-source models (8B–70B) with strong instruction-tuned variants used in our experiments.
+DeepSeek-R1-Distill Guo et al.
+(2025): Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency.
+Other reasoning-enhanced models: Includes S1-32BMuennighoff et al.
+(2025), Limo Ye et al.
+(2025), and FinR1-7B Liu et al.
+(2025b), incorporating reasoning strategies such as financial CoT construction and task-specific fine-tuning.
+Model Name Parameters Enhanced for Reasoning Tasks Context Window Size Close/Open Source Reasoning Enhanced Training Data
 GPT-4o Unknown No 128k Closed -
-GPT-o1 Unknown Yes 128k Closed Public and properity data (Human-annotated CoT, MCTS-assisted Synthetic data)
-GPT-o3-mini Unknown Yes 128k Closed Public and properity data
+GPT-o1 Unknown Yes 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
+GPT-o3-mini Unknown Yes 128k Closed Public and proprietary data
 DeepSeek-V3 671B No 128k Open -
 DeepSeek-R1 671B Yes 128k Open Cold-start data generation, post-processing data
+GPT-4.5 Unknown No 128k Closed -
 Qwen2.5-72B-Instruct 72B No 128k Open -
@@
 Llama3.3-70B-Instruct 70B No 128k Open -
+Qwen2.5-32B-Instruct 32B No 128k Open -
 DeepSeek-R1-Distill-Qwen-32B 32B Yes 128k Open Distilled from R1
+Limo 32B Yes 128k Open Public and proprietary data with Qwen filteration
+S1-32B 32B Yes 128k Open Public and proprietary data + original data, filtered by quality, difficulty, diversity
+Qwen/QwQ-32B 32B Yes 128k Open -
+Qwen2.5-14B-Instruct 14B No 128k Open -
 DeepSeek-R1-Distill-Qwen-14B 14B Yes 128k Open Distilled from R1
@@
 Llama3.1-8B-Instruct 8B No 128k Open -
+Qwen2.5-7B-Instruct 7B No 128k Open -
+FinR1-7B 7B Yes 128k Open 60k CoT data for financial tasks
 Table 2: Summary of evaluated large language models, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies.
@@
 2.3 Evaluation Settings
-In each dataset, the LLMs are presented with a context consisting of financial reports, financial tables, equations, and mathematical symbols, along with a question.
-Insired by previous work Zhao et al.
-(2024), a generic prompt is used to instruct the model to answer the question based on the provided context.
-The models are required to extract relevant information and perform mathematical computations using numbers and concepts derived from the context.
-This necessitates strong logical reasoning abilities within financial contexts.
-The results are typically presented in a mathematical format, but due to variations in numerical representation—such as percentages and different rounding strategies—the final outputs can differ.
-Furthermore, reasoning-focused models, such as DeepSeek-R1 and GPT-o1, tend to provide a step-by-step reasoning process before arriving at the final answer.
-Inspired by DocMath, we employ an LLM-as-judge approach to evaluate model performance.
-First, an LLM (GPT-3.5-Turbo) is used to extract the final answer from the model’s output.
-Then, a mathematical comparison method is applied to determine whether the result is correct.
-For large models like DeepSeek-V3 and DeepSeek-R1, or closed-source models like the GPT series, we conduct our experiments using APIs provided by togetherAI [^4] https://www.together.ai and OpenAI [^5] https://platform.openai.com.
-Specifically, we utilize GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06 for our evaluations.
-For other models with 72B parameters or fewer, we generate results using the VLLM framework on 4 A100 GPUs (80GB memory per GPU), with a maximum generation length of 512 tokens.
-2.4 Fino1: LLMs for Financial Reasoning
-To gain a deeper insight into how domain-specific reasoning enhancement benefits financial tasks compared to general reasoning enhancement, we developed Fino1, a reasoning-enhanced LLM specifically designed for financial tasks.
-WTo enhance the domain-specific reasoning abilities of LLMs, we first construct high-quality training data with domain-specific reasoning paths based on the FinQA dataset.
-FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables.
-This comprehensive dataset integrates both textual and tabular financial data, presenting a moderate level of difficulty, making it well-suited for training reasoning-based models.
-For reasoning path generation, we utilize GPT-4o, as it demonstrates consistently strong performance across multiple datasets (as reported in Table 3), achieving results comparable to the highest-performing models on FinQA.
-Followed by HuatuoGPT-o1 Chen et al.
-(2024), we adopt a similar strategy to refine the CoT reasoning paths generated by GPT-4o.
-The refinement process involves the use of a verifier to assess the correctness of the generated CoT.
-If the verifier rejects the reasoning path, the model iteratively refines it by applying a strategy sampled from one of the following: (1) Backtracking – Revisiting previous steps to correct logical inconsistencies.
-(2) Exploring New Paths – Generating alternative reasoning routes when the current path fails.
-(3) Verification – Cross-checking intermediate steps to ensure logical soundness.
-(4) Correction – Making targeted adjustments to rectify errors.
-The CoT refinement process begins with an initial prompt to generate a reasoning path and an answer.
-The verifier then evaluates the response.
-If the answer is incorrect, one of the above strategies is applied iteratively to refine the CoT process until a correct and logically sound answer is generated.
-2.4.1 Mode Training
-After curating the reasoning data, we fine-tune Llama3.1-8B-Instruct, selected as the optimal backbone model based on the performance of 8B models.
-First, the curated dataset is used for supervised fine-tuning (SFT), where the question serves as input, and the reasoning path along with the final answer serves as output, enabling the model to learn structured financial reasoning.
-Next, we apply reinforcement learning with a reward model trained to assess the quality of generated reasoning paths.
-Inspired by prior work, we employ a verifier-based sparse reward mechanism and optimize the model using Proximal Policy Optimization (PPO).
-The verifier evaluates the correctness and coherence of the reasoning path; if an error is detected, the model iteratively refines the reasoning steps through backtracking, exploring alternative paths, verification, and correction.
-This two-stage approach enhances logical consistency, financial reasoning depth, and equation-solving accuracy, ultimately improving model performance across financial benchmarks.
-During the SFT phase, the model was trained with a learning rate of 5e-6 for three epochs, using a maximum sequence length of 8192 tokens.
-LoRA (Low-Rank Adaptation) was applied to update parameters efficiently, reducing computational costs while maintaining performance.
-In the subsequent RL stage, the learning rate was lowered to 5e-7, and the model was trained for an additional three epochs.
-The reward model, previously trained in Chen et al.
-(2024), was employed to evaluate the quality of the model’s reasoning paths.
-This reward signal guided optimization through Proximal Policy Optimization (PPO), refining the model’s ability to generate accurate and coherent financial reasoning.
+In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question.
+Following prior work Zhao et al.
+(2024a), we use a generic prompt (see appendix D for more details) to instruct models to answer based on the given context.
+Models must extract relevant information and perform mathematical reasoning using numbers and concepts from the context.
+While outputs are typically in mathematical form, variations in formats (e.g., percentages, rounding) may cause differences.
+Reasoning-focused models like DeepSeek-R1 and GPT-o1 often generate step-by-step reasoning.
+Inspired by DocMath Zhao et al.
+(2024a), we adopt an LLM-as-judge evaluation approach: GPT-3.5-Turbo extracts final answers, followed by a mathematical comparison to assess correctness.
+For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series), we run evaluations via APIs from TogetherAI [^2] https://www.together.ai and OpenAI [^3] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06.
+For models up to 72B parameters, we use the VLLM framework on 8 A100 GPUs (80GB each), with a max generation length of 512 tokens.
+2.3.1 Reasoning Data Curation
+To enhance domain-specific reasoning capabilities in LLMs, we curate a high-quality training dataset by integrating diverse financial reasoning benchmarks.
+Our dataset combines a wide range of reasoning tasks—covering numerical computation, table-text integration, cross-table correlation, and long-context understanding—reflecting the complexity of real-world financial documents.
+Specifically, we incorporate data from several well-established benchmarks, including FinQA Chen et al.
+(2022), DocFinQA Reddy et al.
+(2024b), TAT-QA Zhu et al.
+(2021), DocMath-Eval Zhao et al.
+(2024b), and BizBench-QA Krumdick et al.
+(2024).
+These datasets collectively provide rich and diverse coverage across task types and difficulty levels, supporting the development of robust financial reasoning models.
+FinQA provides QA pairs combining textual and tabular financial data; DocFinQA extends this to long-context documents; TAT-QA focuses on complex numerical reasoning; DocMath-Eval (CompLong and SimpShort subsets) targets reasoning over multi-table and single-table scenarios; and BizBench-QA enriches the corpus with code-based financial reasoning tasks.
+Please see Appendix B for detailed descriptions of each dataset.
+2.4 Data Collection and Reasoning Path Construction
+To ensure a balanced dataset of both short and long context samples, we collect all data from the FinQA train set, TATQA, DocMath-CompLong, and DocMath-SimpShort.
+Additionally, we sample 1,000 examples each from BizBench-QA and DocFinQA to balance short and long reasoning tasks.
+In total, we gather 9,851 financial QA samples for reasoning path construction.
+Inspired by Ye et al.
+(2025); Muennighoff et al.
+(2025), which suggest that simple QA pairs may not effectively improve reasoning capabilities, we apply a filtering strategy to retain more challenging examples.
+Specifically, we use Qwen-14B-Instruct to generate answers for each question and employ GPT-3.5-Turbo as a verifier.
+QA pairs correctly answered by Qwen-14B are removed, resulting in a more difficult training subset of 3,472 QA pairs for optimizing more advanced models.
+2.5 Reasoning Path Generation and Refinement
+For reasoning path generation, we use GPT-4o, chosen for its strong performance across datasets (Table 3) and its stability in generating high-quality outputs.
+Following the CoT refinement framework from HuatuoGPT-o1 Chen et al.
+(2024), we employ a verifier-based iterative process to refine the reasoning paths generated by GPT-4o.
+If a reasoning path is rejected by the verifier, refinement is guided by one of the following strategies: (1) Backtracking: Revisiting previous steps to correct logical inconsistencies.
+(2) Exploring New Paths: Generating alternative reasoning routes when the current path fails.
+(3) Verification: Cross-checking intermediate steps to ensure logical soundness.
+(4) Correction: Making targeted adjustments to rectify specific errors.
+The CoT refinement process begins with an initial prompt to generate a reasoning path and final answer.
+If the verifier identifies the answer as incorrect, one of the above strategies is applied iteratively until a correct and logically consistent answer is produced.
+2.5.1 Model Training
+We develop our financial reasoning models based on curated reasoning paths, focusing on smaller-sized models for efficient real-world deployment.
+We select LLaMA-3.1-8B-Instruct and Qwen2.5-14B-Instruct as backbone models, given their strong performance among small and mid-sized LLMs.
+We adopt a two-stage training approach.
+In the first stage, we perform Supervised Fine-Tuning (SFT), using questions as input and reasoning paths with final answers as output.
+For the 8B model, we use only FinQA reasoning paths to assess the impact of domain-specific fine-tuning.
+For the 14B model, we use the full set of sampled and filtered data to leverage its greater capacity for diverse and complex reasoning.
+In the second stage, we apply RL using a verifier-based sparse reward mechanism and optimize with Proximal Policy Optimization (PPO), following Chen et al.
+(2024).
+A reward model evaluates the correctness and coherence of the reasoning path.
+If errors are found, the model iteratively refines its outputs through backtracking, alternative path exploration, verification, and correction, improving logical consistency and equation-solving accuracy.
+During SFT, models are trained for three epochs with a learning rate of 5e-6 and a maximum sequence length of 8192 tokens.
+LoRA is applied for efficient parameter updates.
+In the RL stage, training continues for three epochs with a reduced learning rate of 5e-7, guided by the reward model to refine reasoning quality via PPO.
 3 Results
-3.1 Overall Results Analysis
-The resutls of all these models on three dataset are listed in Table 3.
-Among all models, DeepSeek-R1 achieves the highest overall performance (68.93), driven by its strong XBRL-Math results, with its distilled variants, DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B, closely following.
-Among large-scale models, GPT-4o performs well across datasets but is outperformed by DeepSeek-R1 due to its lower XBRL-Math score.
-General-purpose 70B models, such as Qwen2.5-72B-Instruct and Llama3.3-70B-Instruct, achieve strong results, with Llama3.3-70B showing significant improvement over its predecessors.
-Surprisingly, these models also surpass the reasoning-focused GPT-o1 and approach GPT-o3-mini, suggesting that general-purpose reasoning strategies do not always translate effectively to financial tasks.
-In the 8B category, Fino1-8B (61.03) outperforms other models, demonstrating robust task optimization.
-Notably, reasoning-enhanced models exhibit mixed results: while they improve XBRL-Math scores, they tend to perform worse on FinQA and DM-Simplong.
-This suggests that different reasoning strategies yield varying benefits across financial reasoning tasks, emphasizing the importance of task-specific optimization.
-Models FinQA DM-Simplong XBRL-Math Average
-GPT-4o 72.49 60.00 72.22 68.24
-GPT-o1 49.07 56.00 74.44 59.84
-GPT-o3-mini 60.87 59.00 76.67 65.51
-DeepSeek-V3 73.20 53.00 76.67 67.62
-DeepSeek-R1 65.13 53.00 86.67 68.93
-Qwen2.5-72B-Instruct 73.38 59.00 67.78 66.72
-Qwen2.5-72B-Instruct-Math 69.74 42.00 83.33 65.69
-DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 68.80
-Llama3-70B-Instruct 58.92 41.00 56.67 52.20
-Llama3.1-70B-Instruct 63.18 48.00 63.33 58.17
-Llama3.3-70B-Instruct 68.15 54.00 70.00 64.05
-DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 68.97
-DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 63.90
-DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 53.36
-Llama3-8B-Instruct 41.97 29.00 48.89 39.95
-Llama3.1-8B-Instruct 54.13 34.00 62.22 50.12
-Fino1-8B 60.87 40.00 82.22 61.03
-Table 3: Performance of different LLMs on three tested financial datasets.
-From our results, we identify several key insights regarding the performance of large language models (LLMs) in financial reasoning tasks:
-(1) General reasoning enhancements do not necessarily translate to improvements in financial tasks.
-While models fine-tuned for general reasoning, such as GPT-o1 and DeepSeek-R1, demonstrate strong performance in mathematical and logical reasoning tasks, they do not consistently outperform other models in financial benchmarks.
-For example, GPT-o1 scores 49.07 in FinQA and 56.0 in DM-Simplong, trailing behind models like GPT-4o that are not explicitly optimized for general reasoning.
-Also, DeepSeek-R1 also under performan V3 in these tasks.
-This suggests that financial reasoning requires more than just enhanced logical and mathematical capabilities; it demands a deep understanding of domain-specific terminology, structured financial statements, and numerical patterns.
-One possible explanation for this phenomenon is that reasoning-focused fine-tuning, which emphasizes general problem-solving strategies, may inadvertently weaken a model’s retention of financial knowledge.
-This is particularly relevant for tasks requiring precise comprehension of financial terms, balance sheet structures, and industry-specific metrics, where models without targeted financial fine-tuning may struggle.
-Additionally, reasoning-enhanced models might prioritize general logical inference over domain-specific factual recall, which could lead to a decrease in performance when interpreting specialized financial contexts.
-These findings indicate that while general reasoning improvements enhance broad problem-solving skills, they do not necessarily translate into domain-specific advantages, highlighting the need for specialized financial training in LLMs.
-(2) Different reasoning enhancement strategies yield varying improvements depending on the financial reasoning task format.
-By comparing the performance of GPT-4o vs.
-GPT-o1, Qwen2.5-72B-Instruct vs.
-Qwen2.5-72B-Instruct-Math, and Llama models vs.
-the DeepSeek-R1 distilled Llama models, we observe that reasoning-enhanced models do not always improve financial task performance.
-GPT-o1 underperforms GPT-4o in FinQA (49.07 vs.
+Table 3 presents the performance of various LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong.
+Our models, Fino1-8B and Fino1-14B, demonstrate strong and consistent results across tasks, outperforming several larger and closed-source models.
+Fino1-14B achieves an average score of 60.25, surpassing many larger-scale models such as LLaMA-3.3-70B-Instruct (56.04), Qwen2.5-72B-Instruct (53.71), GPT-o1-preview (54.05), GPT-o3-mini (57.89), and even GPT-4.5 (60.43).
+It also achieves the best performance among all models on two benchmarks—DM-Simplong (60.00) and XBRL-Math (86.67).
+These results demonstrate that high-quality, domain-specific reasoning training can enable smaller models to rival—and even surpass—the performance of significantly larger and proprietary LLMs.
+Fino1-8B also performs competitively with an average score of 50.77, surpassing other 8B models and even some 14B and 32B models.
+Both models excel on XBRL-Math, where Fino1-14B achieves the top score of 86.67.
+Moreover, Fino1-14B and Fino1-8B significantly outperform FinR1-7B, a financial reasoning-enhanced model trained on a much larger 60K task-aligned dataset, demonstrating that larger data volume does not necessarily lead to better performance, and that carefully curated, high-quality reasoning data—even in smaller quantities—can yield superior results.
+Models FinQA DM-Simplong XBRL-Math DM-Complong Average
+GPT-4o 72.49 60.00 72.22 39.33 61.01
+GPT-o1-preview 49.07 56.00 74.44 36.67 54.05
+GPT-o3-mini 60.87 59.00 76.67 35.00 57.89
+DeepSeek-V3 73.20 53.00 76.67 42.33 61.30
+DeepSeek-R1 65.13 53.00 86.67 38.67 60.87
+GPT-4.5 68.94 59.00 74.44 39.33 60.43
+Meta-Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
+Llama-3.1-70B-Instruct 63.18 48.00 63.33 34.33 52.21
+Llama-3.3-70B-Instruct 68.15 54.00 70.00 32.00 56.04
+Qwen2.5-72B-Instruct 73.38 59.00 67.78 14.67 53.71
+Qwen2.5-Math-72B-Instruct 69.74 42.00 83.33 5.00 50.02
+DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 30.67 59.27
+Qwen2.5-32B-Instruct 73.11 56.00 65.56 30.00 56.17
+Qwen/QwQ-32B 61.22 46.00 84.44 20.00 52.92
+DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 24.67 57.40
+Limo 63.44 45.00 61.11 15.33 46.22
+S1-32B 66.81 53.00 84.44 24.00 57.06
+Qwen2.5-14B-Instruct 67.44 59.00 57.78 26.67 52.72
+DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 21.00 53.18
+DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 15.67 43.94
+Meta-Llama-3-8B-Instruct 41.97 29.00 48.89 6.00 31.47
+Llama-3.1-8B-Instruct 54.13 34.00 62.22 14.30 41.16
+Qwen2.5-7B-Instruct 55.37 41.00 42.22 17.67 39.07
+FinR1-7B 58.74 37.00 30.00 13.67 34.85
+Fino1-8B 60.87 40.00 82.22 20.00 50.77
+Fino1-14B 70.01 60.00 86.67 24.33 60.25
+Table 3: Performance of different LLMs on four financial reasoning datasets.
+Beyond overall performance, the comparison across models provides additional insights that can guide the development of LLMs for financial applications:
+(1) General reasoning enhancements do not necessarily improve financial task performance.
+Models like GPT-o1 and DeepSeek-R1, though strong in general reasoning, underperform on financial benchmarks, highlighting the importance of domain-specific knowledge and structured numerical reasoning.
+Moreover, reasoning enhancement strategies based on general-domain data, as seen in S1-32B and Limo, often fail to improve and sometimes even degrade performance compared to their backbone models (e.g., Qwen2.5-32B).
+These results underscore that high-quality, domain-aligned reasoning paths are more effective than general reasoning augmentation, and tailored financial training is crucial for optimal model performance.
+(2) Reasoning enhancements vary in effectiveness across financial tasks.
+Models like GPT-o1 perform worse than GPT-4o on FinQA (49.07 vs.
 72.49) and DM-Simplong (56.00 vs.
-60.00), suggesting that its reasoning strategies, such as chain of thought and reflection-tuning, are optimized for general problem-solving rather than financial-specific logic.
-Similarly, Qwen2.5-72B-Instruct-Math, despite achieving a substantial XBRL-Math boost (67.78 → 83.33), experiences sharp declines in FinQA (69.74 → 42.00) and DM-Simplong (59.00 → 42.00), indicating that mathematical reasoning enhancements help structured financial data but do not generalize well to broader financial tasks.
-In the 8B category, DeepSeek-R1-Distill-Llama-8B (53.36) outperforms Llama3.1-8B-Instruct (50.12) and Llama3-8B-Instruct (39.95), excelling in XBRL-Math (81.11) but showing limited improvements in other financial tasks.
-These results suggest that different reasoning strategies target distinct aspects of financial reasoning.
-DeepSeek-R1 and Qwen2.5-72B-Instruct-Math demonstrate strong performance in symbolic and mathematical reasoning, excelling in structured equation-based tasks.
-However, they do not generalize well to financial question-answering tasks like FinQA and DM-Simplong.
-Similarly, reasoning augmentations in GPT-o1 fail to provide consistent improvements, likely due to the absence of relevant financial reasoning data, highlighting the challenge of adapting broad reasoning techniques to financial logic.
-(3) Scaling model size does not always lead to performance gains in financial tasks.
-While larger models typically achieve better results in standard NLP tasks, their performance in financial tasks does not always scale proportionally with size.
-For instance, DeepSeek-70B (66.73 in FinQA, 53 in DM-Simplong, and 86.67 in XBRL-Math) performs similarly to its smaller distilled variant DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite the latter having less than half the parameters.
-A broader comparison, including Qwen2.5-72B-Instruct, DeepSeek-R1-Distill-Llama-70B, GPT-4o, DeepSeek-V3, and DeepSeek-R1, reveals that performance stabilizes once model size reaches approximately 70B parameters.
-One possible explanation for this trend is the nature of XBRL-Math, which consists of structured financial tables and numerical equations that prioritize precise mathematical computations over complex language-based reasoning.
-In such structured, numerically driven tasks, smaller models can achieve competitive performance by leveraging arithmetic operations and pattern recognition rather than advanced language modeling.
-This suggests that the advantage of larger models diminishes in financial decision-making tasks where structured data comprehension and numerical reasoning play a more crucial role than language-based inference.
-(4) Improving Pre-training Data and Strategies in General Domains Also Benefits Financial Tasks.
-Compared with Llama 3-8B-Instruct, Llama 3.1-8B-Instruct demonstrates a significant performance boost, increasing its average score from 39.95 to 50.12, with substantial improvements in XBRL-Math (62.22 vs.
+60.00).
+Qwen2.5-Math improves in XBRL-Math (67.78 $\rightarrow$ 83.33) but drops sharply in FinQA (69.74 $\rightarrow$ 42.00) and DM-Simplong (59.00 $\rightarrow$ 42.00).
+DeepSeek-R1-Distill-Llama-8B also excels in XBRL-Math (81.11) but shows limited gains elsewhere.
+These results suggest that strategies like CoT or math fine-tuning may offer benefits for structured tasks but often fail to generalize effectively to broader financial reasoning.
+Compared with Qwen2.5-72B-Instruct and its math enhanced version, we can see that while TIR and synthetic building appears to help in equation-heavy tasks like XBRL-Math, its impact on other types of financial tasks is less clear and may be influenced by multiple factors such as task structure, data modality, or domain complexity.
+(3) Larger model size offers limited benefits in financial reasoning.
+While larger models typically excel in general NLP tasks, their gains in financial tasks are not proportional.
+For example, DeepSeek-70B (66.73 FinQA, 53 DM-Simplong, 86.67 XBRL-Math) performs similarly to the smaller DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite having less than half the parameters.
+Across models like Qwen2.5-72B, DeepSeek-R1-Distill-Llama-70B, and GPT-4o, performance tends to plateau around 70B.
+This may reflect the structured nature of tasks like XBRL-Math, where arithmetic precision outweighs complex language reasoning, enabling smaller models to compete effectively.
+(4) Improved pre-training data and strategies in general domains also benefit financial tasks.
+Llama 3.1-8B-Instruct shows a notable performance boost over Llama 3-8B-Instruct (avg.
+50.12 vs.
+39.95), with gains in XBRL-Math (62.22 vs.
 48.89) and FinQA (54.13 vs.
 41.97).
-Similarly, among the 70B-Instruct models, Llama 3.1-70B-Instruct outperforms Llama 3-70B-Instruct (58.17 vs.
-52.20), with the largest improvement in DM-Simplong (48.00 vs.
-41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05 vs.
-58.17), excelling in FinQA (68.15 vs.
-63.18) and XBRL-Math (70.00 vs.
-63.33), highlighting the benefits of dataset expansion and refined training strategies.
-More specifically, Llama 3.1 incorporates a more extensive dataset, covering over 15 trillion tokens across multiple languages and domains, leading to notable performance gains.
-Llama 3.3 further benefits from advancements in post-training techniques, such as online preference optimization, which enhances core performance while significantly reducing training and inference costs.
-These refined optimization strategies in pre-training not only improve general capabilities but also enhance performance in domain-specific tasks.
-(5) Existing reasoning models struggle with long financial contexts and cross-table correlations.
-Many financial reasoning tasks require models to extract and synthesize information from lengthy reports, structured tables, and numerical equations.
-However, performance discrepancies across datasets suggest that current LLMs struggle with this challenge.
-For example, models that perform well in FinQA, which primarily involves textual financial reasoning, do not necessarily excel in DM-Simplong, which integrates numerical and symbolic information.
-GPT-4o achieves 72.49 in FinQA but drops to 60.0 in DM-Simplong, indicating difficulties in integrating mathematical symbols with textual reasoning.
-Furthermore, models that score highly in XBRL-Math, such as DeepSeek-R1-Distill-Qwen-32B (84.44) and DeepSeek-R1 (86.67), highlight a growing but incomplete capability to process structured financial tables.
-Notably, all reasoning-enhanced models fail to show improvements in handling long financial contexts and multi-table correlations.
-This is particularly concerning given that financial reports are inherently long and include complex interlinked tables.
-Though smaller models perform well in structured inputs like XBRL-Math, their performance drops significantly—by nearly 30%—when handling long contexts.
-For example, DeepSeek-R1-Distill-Llama-8B achieves 81.11 in XBRL-Math but drops to 33.0 in DM-Simplong.
-This suggests that while smaller models can recognize structured patterns, they struggle with long-range dependencies and multi-step reasoning.
-Addressing these limitations requires improvements in long-context retention, structured data comprehension, and multi-modal document understanding to enhance financial reasoning capabilities.
-(6) Reasoning enhancement basd on financial data improve performance of all datasets.
-Compared with Llama 3.1-8B-Instruct, DeepSeek-R1-Distill-Llama-8B, which incorporates reasoning enhancement data distilled from DeepSeek-R1, shows only a modest improvement (50.12 to 53.36).
-However, our model, Fino1-8B, trained with reasoning paths distilled from GPT-4o using a backtracking strategy to iteratively refine the reasoning process, achieves a substantial performance boost across all datasets, with an average score of 61.03—a 10.91-point increase over Llama 3.1-8B-Instruct.
-Notably, Fino1-8B outperforms its counterparts in FinQA (60.87 vs.
+Similarly, Llama 3.1-70B-Instruct outperforms Llama 3-70B (58.17 vs.
+52.20), especially in DM-Simplong (48.00 vs.
+41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05), driven by improvements in FinQA (68.15) and XBRL-Math (70.00).
+These gains stem from larger, more diverse training corpora (e.g., 15T+ tokens in Llama 3.1) and enhanced post-training techniques like online preference optimization.
+Such strategies improve general reasoning while also boosting performance in domain-specific tasks like financial reasoning.
+(5) Existing reasoning models struggle with long contexts and cross-table correlations.
+Financial tasks often require integrating information from lengthy reports, structured tables, and equations, yet current models show clear limitations.
+For example, GPT-4o performs well on FinQA (72.49) but drops on DM-Simplong (60.00), indicating challenges in combining textual and symbolic reasoning.
+Models like DeepSeek-R1 and DeepSeek-R1-Distill-Qwen-32B excel in XBRL-Math but fail to generalize.
+Performance on DM-Complong remains low across models (e.g., LLaMA-3.3-70B: 32.00).
+Smaller models like DeepSeek-R1-Distill-Llama-8B also perform poorly with 33.00 for DM-Simplong.
+These results highlight the need for improved long-context modeling and structured data integration in financial reasoning.
+(6) Financial reasoning enhancement improves performance across all tasks.
+Compared to LLaMA 3.1-8B-Instruct (41.16), DeepSeek-R1-Distill-Llama-8B shows a modest gain (43.94).
+In contrast, our Fino1-8B, trained on GPT-4o-distilled reasoning paths with backtracking strategies, achieves a substantial improvement, reaching an average score of 50.77, and consistently outperforms baselines on FinQA (60.87 vs.
 54.13), DM-Simplong (40.00 vs.
-34.00), and particularly in XBRL-Math (82.22 vs.
-62.22), demonstrating the effectiveness of reasoning-driven training.
-It is important to note that our training utilized reasoning paths generated exclusively from FinQA, yet the model exhibited improvements across all datasets.
-This suggests that financial domain-specific reasoning enhancement data can significantly improve models’ comprehensive understanding of financial terminology, long-context comprehension, table-based reasoning, and equation interpretation.
-Furthermore, these results highlight the effectiveness of a combination of reasoning path-building strategies, including backtracking, exploring new paths, verification, and correction, in financial tasks.
-By refining reasoning paths through iterative corrections, these strategies enhance logical consistency, computational accuracy, and overall robustness, demonstrating a promising approach for advancing financial reasoning capabilities in language models.
-Overall, these insights demonstrate that financial reasoning presents unique challenges distinct from general NLP and mathematical reasoning.
-The variability in performance across different tasks underscores the necessity for more financially specialized training methodologies and architectural improvements to enhance financial comprehension, structured data processing, and multi-step numerical reasoning.
-3.2 Error Analysis
+34.00), and XBRL-Math (82.22 vs.
+62.22).
+Fino1-14B further boosts performance, achieving an average score of 60.25, outperforming other 14B models (e.g., Qwen2.5-14B: 52.72) and even rivaling state-of-the-art models like DeepSeek-R1 (60.87) and GPT-4o (61.01).
+Notably, although Fino1-8B and Fino1-14B were trained with limited financial CoTs, both models show improvements across all tasks, highlighting the effectiveness of high-quality reasoning paths and structured training strategies such as backtracking, path exploration, and verification.
+Overall, these insights highlight that financial reasoning poses unique challenges distinct from general NLP and mathematical reasoning.
+The performance variability across tasks underscores the need for specialized training strategies and architectural improvements to better support financial comprehension, structured data processing, and multi-step numerical reasoning.
+4 Conclusion
+This study presents a comprehensive evaluation of 24 general-purpose and reasoning-augmented LLMs on financial reasoning tasks, encompassing financial text understanding, numerical computation, and structured data interpretation.
+Our findings show that existing models, despite strong performance on general reasoning benchmarks, often fail to generalize to financial domains—particularly in handling long contexts, multi-table reasoning, and domain-specific terminology.
+To bridge this gap, we curated a high-quality, diverse financial reasoning dataset and introduced Fino1, a suite of domain-adapted models trained with GPT-4o-generated reasoning paths using chain-of-thought fine-tuning and reinforcement learning.
+Despite being trained on limited data, Fino1 models outperform several larger and advanced LLMs, including GPT-o1, GPT-o3-mini, GPT-4.5, and DeepSeek models (V3 and R1).
+These results highlight the effectiveness of domain-specific reasoning strategies and demonstrate the practical value of tailored model adaptation for real-world financial applications.
+Appendix A Related Work
+Our work builds upon three main research directions: the development of general-purpose and reasoning-enhanced language models, applications of LLMs in finance, and studies on reasoning capabilities in financial tasks.
+A.1 General-Purpose and Reasoning-Enhanced LLMs
+Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning.
+Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting Wei et al.
+(2022), self-reflection Zhang et al.
+(2023), and reinforcement learning from human feedback Ouyang et al.
+(2022).
+These models have shown superior performance in mathematical reasoning Lewkowycz et al.
+(2022), symbolic manipulation Drori et al.
+(2023), and logical inference tasks Chen et al.
+(2023).
+A.2 LLMs in Financial Applications
+The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis Yang et al.
+(2020), market sentiment prediction Yang et al.
+(2023), and automated trading strategies Kou et al.
+(2024).
+Recent work has focused on developing finance-specific models like BloombergGPT Wu et al.
+(2023), FinGPT Yang et al.
+(2023) and PIXIU Xie et al.
+(2023), OpenFinLLMs Xie et al.
+(2024b), which are pretrained on financial corpora to better understand domain-specific terminology and concepts.
+These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
+A.3 Reasoning Capabilities in Financial Tasks
+Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference.
+Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al.
+(2021), mathematical problem-solving in financial contexts Srivastava et al.
+(2024), and structured data interpretation Liu et al.
+(2025a).
+These works highlight the importance of both domain expertise and reasoning abilities in financial applications.
+Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
+Appendix B Details Data Source
+We provide here detailed descriptions of the datasets used to construct our training and evaluation corpus for financial reasoning enhancement.
+  • FinQA Chen et al.
+(2022): FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables.
+It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
+  • DocFinQA Reddy et al.
+(2024b): An extension of FinQA designed for long-context reasoning.
+It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words.
+This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
+  • TAT-QA Zhu et al.
+(2021): A large-scale financial QA dataset created from real financial reports.
+It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison.
+TAT-QA reflects real-world challenges in financial document interpretation.
+  • DocMath-Eval Zhao et al. (2024b): This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
+    • DocMath-CompLong: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
+    • DocMath-SimpShort: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
+  • BizBench-QA Krumdick et al.
+(2024): Contains 14,400 financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode.
+These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
+Figure 1: Error case 1.
+Figure 2: Error case 2.
+Appendix C Error Analysis
 To further investigate why reasoning-enhanced models like DeepSeek-R1 and GPT-o1 underperform in financial tasks compared to general-purpose models such as GPT-4o and DeepSeek-V3, we conduct an error analysis using the FinQA dataset since these reasoning-enhanced models exhibit a significant performance drop on FinQA.
@@
 This highlights a fundamental gap in DeepSeek-R1’s financial reasoning capabilities, particularly in recognizing the accounting conventions used to report share-based compensation.
-Figure 1: Error case 1.
-Figure 2: Error case 2.
-3.3 Future Direction
-Based on the above findings, we observe that developing a powerful reasoning model for financial tasks presents unique challenges distinct from general-domain reasoning.
-Our analysis suggests three key areas for improvement in financial reasoning models:
-(1) Enhancing models’ understanding of financial knowledge and terminology.
-While reasoning-enhanced models demonstrate strong performance in general domains, they often struggle with financial tasks due to a lack of domain-specific knowledge.
-Future work should explore strategies that incorporate financial corpora, structured financial reports, and regulatory documents to improve models’ comprehension of domain-specific terminology and contextual nuances.
-(2) Improving models’ ability to handle multi-table reasoning and long financial contexts.
-Our results highlight that even large models exhibit significant performance drops when processing long financial reports and cross-referencing multiple tables.
-To address this, future reasoning stratigies should be focused on inferring logics among tables or menorization mechanisms to enhance long-context understanding.
-(3) Deeper analysis of reasoning-enhanced strategies in models like GPT-o1 and DeepSeek-R1.
-Although GPT-o1 and DeepSeek-R1 are designed with advanced reasoning enhancements, they fail to show consistent improvements across financial tasks.
-GPT-o1 utilizes "chain-of-thought" prompting, generating intermediate reasoning steps to enhance structured problem-solving.
-However, this reasoning strategy does not lead to improvements in any of the financial benchmarks, as GPT-o1 underperforms compared to GPT-4o across all tasks, scoring 49.07 in FinQA and 56.0 in DM-Simplong.
-In contrast, DeepSeek-R1 introduces self-reflection capabilities through reinforcement learning, enabling the model to iteratively evaluate and refine its reasoning processes without supervised data.
-While this approach helps DeepSeek-R1 achieve strong performance in XBRL-Math (86.67), indicating improved structured numerical reasoning, it also leads to a drop in FinQA (65.13), suggesting challenges in financial text-based reasoning.
-These results highlight a fundamental difference between the two models — GPT-o1’s structured reasoning approach does not translate to domain-specific improvements, whereas DeepSeek-R1’s reinforcement learning strategy enhances structured data processing but remains limited in free form textual financial reasoning.
-Understanding these trade-offs is essential for designing reasoning strategies tailored to financial tasks.
-Developing robust financial reasoning models requires both domain adaptation and architectural advancements to bridge the gap between general reasoning capabilities and financial-specific challenges.
-Future work should focus on integrating financially specialized pretraining, multi-table comprehension techniques, and long-context optimization to build models capable of handling complex financial documents and reasoning tasks effectively.
-4 Related Work
-Our work builds upon three main research directions: the development of general-purpose and reasoning-enhanced language models, applications of LLMs in finance, and studies on reasoning capabilities in financial tasks.
-4.1 General-Purpose and Reasoning-Enhanced LLMs
-Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning.
-Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting Wei et al.
-(2022), self-reflection Zhang et al.
-(2023), and reinforcement learning from human feedback Ouyang et al.
-(2022).
-These models have shown superior performance in mathematical reasoning Lewkowycz et al.
-(2022), symbolic manipulation Drori et al.
-(2023), and logical inference tasks Chen et al.
-(2023).
-4.2 LLMs in Financial Applications
-The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis Yang et al.
-(2020), market sentiment prediction Yang et al.
-(2023), and automated trading strategies Kou et al.
-(2024).
-Recent work has focused on developing finance-specific models like BloombergGPT Wu et al.
-(2023), FinGPT Yang et al.
-(2023) and PIXIU Xie et al.
-(2023), OpenFinLLMs Xie et al.
-(2024b), which are pretrained on financial corpora to better understand domain-specific terminology and concepts.
-These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
-4.3 Reasoning Capabilities in Financial Tasks
-Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference.
-Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al.
-(2021), mathematical problem-solving in financial contexts Srivastava et al.
-(2024), and structured data interpretation Liu et al.
-(2025).
-These works highlight the importance of both domain expertise and reasoning abilities in financial applications.
-Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
-5 Conclusion
-In this study, we evaluated the performance of 16 recent reasoning-augmented LLMs and general-purpose LLMs on financial tasks, analyzing their capabilities in text understanding, numerical reasoning, and structured data processing.
-Our results indicate that while reasoning-enhanced models demonstrate strong performance in general tasks, they often fail to generalize effectively to financial reasoning due to insufficient domain adaptation.
-Additionally, models optimized for structured mathematical reasoning perform well in numerical tasks like XBRL-Math but struggle with financial text comprehension and long-context processing, highlighting the need for financial domain-specific adaptations.
-To address these limitations, we presented Fino1, a reasoning-enhanced LLMs based on Llama-3.1-8B-Instruct, leveraging reasoning paths curated from GPT-4o using only FinQA, a dataset with 5,000 financial reasoning samples.
-Our results demonstrate that, despite being trained on a single dataset, Fino1 achieved an overall performance improvement of approximately 10% across three financial benchmarks.
-This significant improvement underscores the potential of financial domain-specific data in enhancing LLMs’ financial reasoning capabilities.
-Future work can focus on enhancing financial knowledge adaptation, improving multi-table reasoning, and handling long-context financial documents.
-Additionally, refining domain-specific reasoning strategies through structured data reasoning, retrieval-augmented approaches, and domain-aware pretraining will be critical for bridging the gap between general reasoning capabilities and financial-specific challenges.
 Limitations
@@
 Fourth, our reasoning path construction approach is simplified, as we generate paths using a single method (GPT-4o), whereas exploring multiple reasoning path generation strategies—such as ensemble approaches or human-annotated paths—could lead to more robust financial reasoning capabilities.
+C.1 Future Direction
+Developing effective reasoning models for financial tasks poses distinct challenges compared to general-domain reasoning.
+Our findings suggest three key directions for improvement: (1) Enhancing financial knowledge and terminology understanding through domain-specific pretraining using financial corpora, structured reports, and regulatory documents; (2) Improving multi-table reasoning and long-context comprehension by developing strategies that better infer logic across tables and optimize memory mechanisms for processing lengthy financial documents; and (3) Refining reasoning-enhanced strategies, as current approaches like GPT-o1’s chain-of-thought prompting and DeepSeek-R1’s self-reflection via reinforcement learning yield mixed results—highlighting the need for more tailored reasoning strategies to address both structured numerical and unstructured textual financial reasoning.
 Ethics Statement
@@
 As with all AI-driven financial models, careful human oversight remains essential, and future work must focus on enhancing factual consistency, mitigating biases, improving domain adaptation, and ensuring alignment with regulatory frameworks before such models can be considered for real-world financial deployment.
+Appendix D Prompt
+Prompt Template for DocMath-CompLong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context.
+You need to first think through the problem step by step, documenting each necessary step.
+Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’.
+The final answer should be a numeric value.
+###Context <Full document here> ### Input <the question based on the document> Lets think step by step to answer the given question.
+### Output
+Prompt Template for DocMath-Simplong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context.
+You need to first think through the problem step by step, documenting each necessary step.
+Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’.
+The final answer should be a numeric value.
+###Context <simple long context here> ### Input <the question based on the context> Lets think step by step to answer the given question.
+### Output
+Prompt Template for FinQA task Please answer the given financial question based on the context. Context: <context> Question: <the question based on the context> Answer:
+Prompt Template for XBRL-Math task You are a financial expert tasked with carefully reading, analyzing, and answering the following eXtensible Business Reporting Language.
+Please follow the steps below: INPUT: Read the eXtensible Business Reporting Language (XBRL) question: <XBRL related question>, formula: <the formula related to this question>, and the explanation: <term explanation>.
+Provide only the final answer which is the numerical result of the calculation.
+For formulas like ROI, provide percentages.
+Never use the percent symbol in percentages.
+OUTPUT:

### v3

@@
-Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance
+Fin-o1: On the Transferability of Reasoning-Enhanced LLMs and Reinforcement Learning to Finance
 Abstract
-While large language models (LLMs) have shown strong general reasoning capabilities, their effectiveness in financial reasoning—which is crucial for real-world financial applications—remains underexplored.
-In this study, we conduct a comprehensive evaluation of 24 state-of-the-art general and reasoning-focused LLMs across four complex financial reasoning tasks involving financial text, tabular data, and equations.
-We assess key capabilities such as numerical reasoning, tabular interpretation, financial terminology comprehension, long-context understanding, and equation-based problem solving.
-Our analysis reveals that while data quality and pretraining contribute to performance, general techniques like chain-of-thought (CoT) fine-tuning offer limited gains in financial tasks.
-To address this, we propose two domain-adapted models, Fino1-8B and Fino1-14B, trained with CoT fine-tuning and reinforcement learning using domain-specific reasoning paths.
-Our models are trained on a carefully curated dataset integrating high-quality examples from diverse sources, covering financial reports, tables, equations, and structured XBRL texts.
-Despite limited training data, they achieve an 7–9% performance improvement, outperforming several advanced LLMs—including GPT-o1, GPT-o3-mini, GPT-4.5, and comparable with DeepSeek models (V3 and R1)—demonstrating strong practical value in resource-constrained scenarios.
-Our findings highlight the need for domain-specific adaptations in financial reasoning, and we release all datasets, models, and code for future research [^1] https://github.com/The-FinAI/Fino1.
+As the fundamental capability behind decision-making in finance, financial reasoning poses distinct challenges for LLMs.
+Although reinforcement learning (RL) have boosted generic reasoning, the progress in finance is hindered by the absence of empirical study of building effective financial chain-of-thought (CoT) corpus, a systematic comparison of different RL methods, and comprehensive benchmarks.
+To address these gaps, we introduce FinCoT, the first open high-fidelity CoT corpus for finance, distilled from seven QA datasets by a novel three-stage pipeline that incorporates domain supervision, iterative LLM refinement, and difficulty-aware filtering.
+Based on FinCoT, we develop Fin-o1, the first open financial reasoning models trained via supervised fine-tuning and GRPO-based RL.
+Our models outperform existing financial reasoning models and SOTA general models such as GPT-o1, DeepSeek-R1, and GPT-4.5.
+We also investigate the effectiveness of three different RL methods in improving domain-specific reasoning, offering the first such empirical study.
+We finally propose FinReason, the first financial reasoning benchmark covering multi-table analysis, long-context reasoning, and equation-based tasks, and evaluate 29 LLMs.
+Our extensive experiments reveal general reasoning models excel on standard benchmarks yet exhibit obvious performance degradation in financial contexts; even finance-tuned models like Dianjin-R1 and FinR1 degrade on lengthy documents.
+In contrast, our Fin-o1 models consistently outperform their backbones and larger GPT-o1 and DeepSeek-R1, confirming the effectiveness of our data building and model training strategy.
+Our study further shows that GRPO yields reliable gains whereas PPO and DPO do not, highlighting the need for targeted data and optimisation rather than scale alone.
+These findings show that financial reasoning has unique challenges unmet by mere model scaling or generic techniques, highlighting the need for specialized data, models, and optimization.
+We release all datasets, models, and codes to support future research.
+111 Dataset: https://huggingface.co/datasets/TheFinAI/FinCoT
+      Fin-o1: https://huggingface.co/TheFinAI/Fin-o1-8B
+                https://huggingface.co/TheFinAI/Fin-o1-14B
+      Code: https://github.com/The-FinAI/Fino1
 1 Introduction
-Advancements in large language models (LLMs) have demonstrated remarkable performance across various natural language pocessing tasks, including content generationVan Veen et al.
-(2024); Zhang et al.
-(2024), language translation Xu et al.
-(2024), and sentiment analysis Xing (2024); Miah et al.
-(2024).
-More recently, reasoning-enhanced models such as OpenAI’s o1 and DeepSeek’s R1 have been developed to extend LLM capabilities in complex reasoning Guo et al.
-(2025); Jaech et al.
-(2024).
-These models exhibit significant improvements, particularly in mathematical and logical tasks that require complex reasoning Temsah et al.
-(2024); Zhong et al.
-(2024).
-However, despite their success in general reasoning, their performance in the financial domain remains largely unexplored.
-Financial reasoning is essential in the financial domain, where professionals must extract key insights from lengthy reports and multi-table documents.
-It requires not only numerical computation but also a deep understanding of financial terminology, regulations, and economic concepts Xie et al.
-(2023, 2024a).
-Unlike general reasoning tasks, financial reasoning involves (1) domain-specific language comprehension, (2) reasoning over financial concepts and numbers, and (3) interpreting structured data such as financial tables Xie et al.
-(2024b).
-Evaluating LLMs in this context is critical for closing the gap between general AI reasoning and real-world financial applications.
-To address this gap, we firstly comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks, to provide insights into the strengths and limitations of existing reasoning models in financial applications.
-We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks.
-(2) Impact across financial tasks – Analyzing how the reasoning enhancement process influences LLM performance across different types of financial tasks.
-(3) Scaling and performance gap – Investigating the differences between smaller reasoning models and large-scale LLMs in financial reasoning.
-(4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
-We evaluated 24 LLMs from the GPT Achiam et al.
-(2023), LLaMA Dubey et al.
-(2024), DeepSeek Liu et al.
-(2024), and Qwen Yang et al.
-(2024) families on four financial reasoning benchmarks: FinQA Chen et al.
-(2021), DM-Simplong Zhao et al.
-(2024a), XBRL-Math Wang et al.
-(2025) and DM-Complong.
-These datasets span financial texts, tables, and equations, enabling evaluation of models in numerical reasoning, domain-specific terminology comprehension, long-context processing, and equation-based reasoning tasks.
-Our evaluation encompasses both general-purpose and reasoning-enhanced models, ranging from smaller-scale models (e.g., LLaMA 3-8B-InstructDubey et al.
-(2024)) to high-capacity models (e.g., GPT-o1Jaech et al.
-(2024)), as well as reasoning-augmented models incorporating strategies such as data sampling and financial Chain-of-Thought (CoT) construction.
-Experimental results reveal three key insights: (1) Reasoning-enhanced models like DeepSeek-R1 and GPT-o1 do not consistently outperform general-purpose models on financial tasks, likely due to challenges related to domain-specific terminology and multi-table reasoning; (2) Performance gains from larger model sizes and improved pretraining tend to plateau beyond 32B or 70B parameters;
-To develop reasoning models tailored for financial domain, we build novel financial reasoning-enhanced large language models Fino1-8B and Fino1-14B based on LLaMA-3.1-8B-InstructDubey et al.
-(2024) and Qwen2.5-14B-InstructYang et al.
-(2024), chosen for their strong performance among smaller-scale LLMs.
-To enhance reasoning capabilities, we apply CoT fine-tuning and reinforcement learning (RL), inspired by the training strategies of HuatuoGPT-o1 Chen et al.
-(2024).
-We leverage a diverse set of high quality financial reasoning datasets—including FinQA Chen et al.
-(2022), TATQA Zhu et al.
-(2021), DocMath (Simpshort and Compshort) Zhao et al.
-(2024b) , DocFinQA Reddy et al.
-(2024a), and BizBench-QA Krumdick et al.
-(2024)—covering tasks such as long-context question answering, complex reasoning over financial documents, and table-text integration.
-We construct high-quality reasoning paths using GPT-4o, combined with a novel data filtering process that selects more challenging samples for training—particularly for the 14B model—ensuring the model learns from difficult, domain-specific reasoning cases rather than generic QA pairs.
-Experimental results show that this targeted training strategy enables our models to surpass the performance of advanced closed-source models such as GPT-o3-mini and GPT-o1, and approach most powerful LLMs like DeepSeek-V3, DeepSeek-R1, and GPT-4o, demonstrating that domain-specific reasoning-enhanced training with limited, high-quality data can be highly effective.
-To the best of our knowledge, this work presents the first comprehensive study of reasoning-enhanced LLMs for financial tasks involving diverse inputs such as financial documents, tables, equations, and structured XBRL texts.
-First, we conduct a systematic evaluation of 24 general-purpose and reasoning-enhanced LLMs across multiple financial reasoning benchmarks.
-Second, we introduce Fino1, the first suite of domain-specific, reasoning-enhanced LLMs, trained using GPT-4o-distilled reasoning paths with a novel data curation and filtering strategy.
-Third, our results show that domain-specific reasoning training with limited high-quality data consistently outperforms general-domain strategies and even surpasses larger closed-source models.
-We release all models, datasets, and code to support future research in financial AI.
-2 Methods
-2.1 Datasets
-To evaluate the financial reasoning capabilities of reasoning-oriented and general LLMs, we select four diverse benchmark tasks.
-(1) FINQA Chen et al.
-(2021) is a large-scale dataset that emphasizes complex numerical reasoning in financial reports, requiring the integration of structured tables and unstructured text, along with a deep understanding of domain-specific terminology.
-(2) DocMath (Simplong, Complong) Zhao et al.
-(2024a) evaluates LLMs’ numerical reasoning abilities over long, specialized documents with multi-tiered tables, testing their capacity to extract and reason over extended contexts.
-The simplong subset contains relatively easier questions, while the complong subset includes more complex questions that require advanced reasoning across tables and context.
-(3) XBRL-Math Wang et al.
-(2025) targets reasoning over structured financial data in XBRL filings, which involve standardized taxonomies, hierarchical disclosures, and formulas, challenging models to interpret intricate numerical relationships.
-Dataset Size Data Types Average Token
-FinQA 1100 Tables and Texts 1,128
-DM-Simplong 100 Tables and Texts 4,330
-DM-Complong 300 Tables and Texts 39,983
-XBRL-Math 90 Texts and Equations 397
-Table 1: Overview of the datasets used in the study.
-More details of the datasets are listed in Table 1.
-By conducting experiments on these three datasets, we aim to evaluate the model’s ability in various aspects, including understanding financial terms, extracting relevant numbers and entities from financial reports from different sources, and reasoning over long contexts and multiple tables.
+Financial reasoning is a critical yet underexplored capability for large language models (LLMs), especially as they are increasingly deployed in high-stakes financial applications.
+Unlike general reasoning tasks, such as mathematical problem solving or code generation, financial reasoning requires models to interpret domain-specific language, analyze structured data like financial tables xie2024open , and extract insights from lengthy, complex documents reddy2024docfinqa .
+It demands not only numerical computation, but also contextual understanding of financial regulations, economic concepts, and ambiguous textual information xie2023pixiu ; xie2024finben .
+Recent advances in reinforcement learning (RL) methods, including proximal policy optimization (PPO) schulman2017proximal , direct preference optimization (DPO) rafailov2023direct , and generalized reinforcement preference optimization (GRPO) tang2024generalized , have substantially enhanced the reasoning capabilities of large language models.
+These RL-based approaches form the foundation of state-of-the-art (SOTA) reasoning models such as OpenAI’s o1 jaech2024openai and DeepSeek’s R1 guo2025deepseek , achieving notable improvements on mathematical and logical problem-solving benchmarks temsah2024openai ; zhong2024evaluation .
+By directly optimizing for reasoning performance, RL strategies have become essential for advancing the complex reasoning abilities of LLMs.
+However, improving LLMs with RL for financial reasoning is yet hindered by three critical challenges (As shown in Table 1).
+First, the absence of high-quality, publicly available chain-of-thought (CoT) datasets for financial reasoning limits the ability to model domain-specific reasoning processes.
+Existing resources are either proprietary jaech2024openai or adapted from general domains guo2025deepseek ; yang2024qwen2 , failing to capture the nuances of financial analysis.
+No empirical study has asked which source documents or path-generation strategies actually produce effective financial CoT traces.
+Second, while RL methods have enhanced general reasoning, their effectiveness in financial tasks remains unexplored, with no open-source models demonstrating domain-adapted improvements.
+The impact of different RL algorithms on financial reasoning remains untested: RL methods like PPO, DPO and GRPO have never been compared side-by-side, leaving practitioners without guidance on the best optimiser for the domain.
+Third, there is no comprehensive benchmark to systematically evaluate financial reasoning across key challenges such as domain-specific language understanding xie2024open , multi-table reasoning zhao2024docmath , long-context document analysis reddy2024docfinqa , and equation-based reasoning ma2023xbrl .
+Existing benchmarks like PIXIU xie2023pixiu , FinBen xie2024finben , and BizBench krumdick2024bizbench address only parts of this space.
+Moreover, current studies only evaluated a few LLM’s financial reasoning ability.
+Table 1: Comparison of Existing Reasoning LLMs and Financial Benchmarks.
+SFT indicates super- vised fine-tuning.
+RL specifies the reinforcement learning method used.
+CoT data is categorized into Guide, indicate the presence of gold guidelines to direct reasoning generation, IterR means iteratively verification and refinement, DiffF indicates if involving difficulty-aware filtering to emphasize complex cases.
+Evaluation indicates whether the benchmarks involve single-table or multi-table reasoning, long-context analysis, and equation-based reasoning.
+ModN indicates the number of models evaluatd in each work.
+*Indicates works released after Fin-o1.
+Model / Benchmark Domain CoT Data Model Evaluation
+Public Guide IterR CompF SFT RL Single Multi Long Context Equations ModN
+PIXIU xie2023pixiu Financial $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ 9
+FinBen xie2024finben Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ 21
+BizBench krumdick2024bizbench Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ 16
+DocFinQA reddy2024docfinqa Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ 12
+S1-32B muennighoff2025s1 General $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 12
+Limo ye2025limo General $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 6
+HuatuoGPT-o1 chen2024huatuogpt Other $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ PPO $\times$ $\times$ $\times$ $\times$ 10
+DeepSeek-R1 guo2025deepseek General $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ GRPO $\times$ $\times$ $\times$ $\times$ $\times$
+OpenAI-o1 jaech2024openai General $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Fin-R1finr1 * Financial $\times$ $\times$ $\times$ $\times$ $\checkmark$ GRPO $\checkmark$ $\times$ $\times$ $\times$ 10
+DianJin-R1zhu2025dianjin * Financial $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ GRPO $\checkmark$ $\times$ $\times$ $\times$ 12
+Fin-o1 (Ours) Financial $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ GRPO, PPO, DPO $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29
+To address these gaps, we present three key contributions.
+First, we curate FinCoT, the first high-quality, publicly available financial CoT dataset built by a novel three-stage framework that combines domain-aware supervision, iterative LLM refinement, and difficulty-aware filtering.
+We construct FinCoT from diverse sources including FinQA chen2022finqadatasetnumericalreasoning , ConvFinQA chen2022convfinqa , DocFinQA reddy2024docfinqa , TATQA zhu2021tatqaquestionansweringbenchmark , Econ_Logic quan2024econlogicqa , DocMath-Eval zhao2024docmath , and BizBench-QA krumdick2024bizbench .
+Our new framework yields high-fidelity financial CoT data that improves 8B and 14B models beyond SOTA reasoning models such as GPT-o1, DeepSeek-R1, and GPT-4.5.
+Second, we develop Fin-o1 (including Fin-o1-8B and Fin-o1-14) based on FinCoT, the first open-source financial reasoning large language models, trained via a two-phase framework combining supervised fine-tuning and RL-based reasoning enhancement.
+Fin-o1 consistently outperform existing financial reasoning models and SOTA general-reasoning LLMs such as GPT-o1, DeepSeek-R1, GPT-4.5.
+We systematically investigate the effectiveness of PPO schulman2017proximal , DPO rafailov2023direct , and GRPO tang2024generalized in improving domain-specific reasoning, offering the first empirical study on RL methods for financial reasoning.
+Third, we introduce FinReason, the first comprehensive benchmark designed to systematically evaluate the financial reasoning capabilities of LLMs.
+FinReason covers most LLMs with diverse tasks involving domain-specific language, structured data, and long-context documents, providing a robust framework on datasets such as FinQA chen2021finqa , DM-Simplong zhao2024docmath , XBRL-Math wang2025finnlp , and DM-Complong zhao2024docmath .
+We conduct a comprehensive evaluation on FinReason, covering 29 large language models from general-purpose, reasoning-enhanced, and financial-specific families, including GPT, LLaMA, DeepSeek, Qwen, and our Fin-o1 models.
+Our results reveal that general reasoning models, despite their strong performance on generic benchmarks, struggle with financial reasoning tasks that require precise domain understanding, structured data interpretation, and long-context document reasoning.
+Frontier financial reasoning LLMs such as Dianjin-R1 and FinR1 exhibit performance degradation compared to their backbones in long-context tasks, highlighting their limited robustness.
+In contrast, our Fin-o1 models, trained on the FinCoT dataset, consistently outperform their backbone models and surpass larger general-purpose models, demonstrating the importance of domain-specific reasoning supervision.
+Further analysis shows that reinforcement learning methods have varying effectiveness in this context.
+While PPO and DPO fail to yield stable improvements, GRPO achieves consistent gains by aligning optimization with task-specific reasoning objectives.
+These findings illustrate that financial reasoning poses unique challenges that cannot be addressed by scaling model size or applying general reasoning techniques alone, emphasizing the need for specialized data, models, and optimization strategies.
+2 FinCoT Reasoning Training Data
+In this section, we describe the construction of FinCoT, the first high-fidelity dataset of chain-of-thought reasoning paths designed to enhance and evaluate LLMs’ financial reasoning capabilities, as illustrated in Fig.
+1.
+2.1 Data Collection
+To develop a robust chain-of-thought training dataset tailored for financial reasoning, we systematically curate data from diverse financial data sources.
+Distinguished from FinR1 finr1 and Dianjin zhu2025dianjin , which only include FinQA for reasoning tasks while the rest of their datasets, i.e, Finance-Instruct flowers2025finance , FinCorpus duxiaoman2023fincorpus , and others that focus primarily on financial knowledge enhancement [^2] Samples of these data can be found in Appendix B.1, our dataset explicitly provides structured chain-of-thought supervision to enable complex financial reasoning, each contributing distinct reasoning skills [^3] See Appendix B for dataset details.:
+  • FinQA chen2022finqadatasetnumericalreasoning and ConvFinQA chen2022convfinqa : Focused on numerical reasoning over short financial texts and single well-structured tables, supporting the model in basic arithmetic and aggregation tasks.
+  • TATQA zhu2021tatqaquestionansweringbenchmark : Provides semi-structured financial reports with single-table QA tasks, enhancing the model’s ability to reason over less curated, real-world financial tables.
+  • DocMath-Eval zhao2024docmath : Adds mathematical reasoning over structured tables, introducing complex quantitative reasoning patterns.
+  • Econ-Logic quan2024econlogicqa : Includes financial decision-making and logical inference tasks, requiring models to apply domain-specific logic beyond numerical calculation.
+  • BizBench-QA krumdick2024bizbench : Introduces multi-table reasoning tasks, training the model to integrate and synthesize information across multiple data sources.
+  • DocFinQA reddy2024docfinqa : Comprises long-form financial documents with QA tasks, enabling the model to handle extended context and multi-step reasoning over lengthy narratives.
+Figure 1: Overall framework of development of FinCoT, Fin-o1 and FinReason.
+2.2 Iterative Chain-of-Thought Generation and Verification with Domain-Guided Supervision
+Building our raw data collection, we further curate the first dataset containing 9,186 chain-of-thought annotated QA pairs, specifically designed to support SFT and RL for financial reasoning models, via a three-stage CoT generation framework by integrating domain-aware supervision, iterative refinement, and difficulty-aware filtering.
+This framework addresses the limitations of existing chain-of-thought methods, which prompt LLMs such as DeepSeek-R1 or Gemini team2024gemini ; ye2025limo ; muennighoff2025s1 to generate reasoning paths in a single round finr1 ; zhu2025dianjin , sufficient for general tasks but inadequate for financial reasoning which requires precise, multi-step, domain-informed calculations.
+For example, financial tasks like calculating profits require carefully following each calculation step with the right data and rules, where even small mistakes can completely invalidate the final answer.
+In the first stage, we tackle the lack of structured guidance in existing methods by incorporating programmatic steps and expert-curated sub-questions from datasets such as BizBench and ConvFinQA, collected during our data preparation phase.
+While ConvFinQA offers high-quality sub-questions, it does not provide complete question-answer pairs that capture the full financial decision-making process.
+To close this gap, we use GPT-4o to synthesize these sub-questions into combined, holistic questions (see Fig.
+2.
+This allows the model to see both the detailed steps and how they fit into a broader financial reasoning workflow, helping it generate more coherent and logically consistent reasoning paths that mirror domain expert expectations).
+Figure 2: Workflow of curating the combined question and examples of program progress.
+In the second stage, we address the brittleness of single-pass reasoning by introducing an iterative verification and refinement process, enabling the model to correct its reasoning progressively.
+Inspired by HuatuoGPT-o1 chen2024huatuogpt but adapted to the stricter demands of financial reasoning, we prompt GPT-4o to generate step-by-step reasoning paths, which are independently evaluated by a verifier (also GPT-4o).
+Instead of discarding incorrect outputs entirely, our framework forces the model to revisit earlier steps, refine its assumptions, and adjust its reasoning flow based on verifier feedback (see Fig.
+3 in appendix C.).
+This allows the model to detect and fix errors that might otherwise propagate unchecked, ensuring greater accuracy and robustness, especially in complex financial tasks where early missteps can have outsized impacts.
+In the third stage, we introduce a difficulty-aware filtering step to ensure that the final dataset targets genuinely challenging reasoning scenarios (as shown in Fig.
+1).
+Unlike prior methods chen2024huatuogpt ; finr1 that focus only on verifying correctness, we also consider question difficulty, motivated by recent studies showing that easy QA pairs do little to advance reasoning ability ye2025limo ; muennighoff2025s1 .
+We use Llama3.1-8B-Instruct to attempt each question, then apply a verifier (see Section 4.3) to check the answers.
+QA pairs that are correctly solved by the filtering LLM are excluded, which is deemed less challenging.
+3 Fin-o1
+In this section, we propose Fin-o1 (including Fin-o1-8B and Fin-o1-14B) based on FinCoT, the first open source LLMs targeted on financial reasoning tasks.
+Qwen3 are selected as backbone models, given their strong performance among small and mid-sized LLMs (shown in Table 3).
+Following previous studies chen2024huatuogpt ; dubey2024llama , a two-stage training approach, supervised fine-tuning and RL, is adopted to develop the model.
+FinCoT is divided into two parts FinCoT_SFT (7686) and FinCoT_RL (1,500) to support the training.
+Based on FinCoT, we further conduct the first evaluation of different RL methods to assess their impact on financial reasoning tasks.
+3.1 Supervised fine-tuning
+In the first stage, we perform supervised fine-tuning, with the FinCoT_SFT, aiming to guide the model to systematically structure its reasoning process before providing the final answer.
+As shown in the right part of Fig.
+1, each training instance is composed of a prompt, contextual information extracted from textual and tabular data, and a question that requires multi-step reasoning.
+The output target consists of a detailed reasoning path, followed by a summarized answer, emphasizing the importance of step-by-step logical analysis before deriving the final response.
+By structuring the output in this manner, the objective of SFT is to cultivate a disciplined reasoning framework, encouraging the model to first articulate a coherent analytical process, thereby enhancing interpretability and accuracy in complex financial reasoning scenarios.
+After the supervised fine-tuning, we get Fin-o1-SFT.
+3.2 Reinforcement learning
+In the second stage, we conduct a comprehensive evaluation of various RL methods to assess their impact on financial reasoning tasks.
+To effectively evaluate these methods, we specially design targeted reward functions and curate high-quality positive and negative samples tailored to each method.
+This comparative analysis aims to provide insights into the strengths and limitations of each RL method in enhancing reasoning capabilities for complex financial tasks.
+Proximal policy optimization.
+We start from PPO schulman2017proximal .
+It optimizes the policy by maximizing the expected reward while constraining the extent of policy updates, thus preventing excessive divergence from the reference policy.
+Each input $x$ is processed by the Fin-o1-SFT to generate a response with a reasoning path $y^{*}$.
+The reward $r$ is defined based on the verifier feedback when comparing if the $y^{*}$ is consistent with the gold answer $y$.
+To effectively evaluate the generated answer with the gold, GPT-4o is used to verify the correctness of the results.
+$r(x,y,y^{*})=\begin{cases}1&\text{if }\text{verifier}(y,y^{*})=\text{True}\\ 0&\text{otherwise}\end{cases}$ (1)
+Additionally, the reward is adjusted by incorporating the Kullback-Leibler (KL) divergence penalty between the updated policy $\pi_{\theta}$ and the reference policy $\pi_{ref}$ luong2024reft , aiming to prevent excessive deviation from the reference policy while ensuring stable learning.:
+$r^{\prime}(x,y,y^{*})=r(x,y,y^{*})-\beta\cdot KL(\pi_{\theta}\parallel\pi_{ref})$ (2)
+Direct preference optimization.
+DPO rafailov2023direct refines the model’s decision-making by systematically contrasting positive and negative samples through pairwise comparisons.
+In our framework, we extend the conventional DPO setup by implementing a more strategically designed sampling mechanism tailored to the nuances of complex financial reasoning paths.
+The iterative CoT reasoning path generation process is leveraged in this step to effectively curate high-quality positive-negative samples.
+As detailed in 2.2, this process involves employing GPT-4o to generate reasoning paths that undergo iterative verification and correction.
+Rather than employing low-quality negative samples that diverge significantly from the positive samples, we implement a targeted selection strategy that designates the last failed reasoning path $R_{n-1}$ as the negative sample.
+This path is specifically chosen as it closely mirrors the input query in logical structure while containing identifiable reasoning flaws, making it a more informative negative sample.
+By focusing on plausibility and logical coherence, our sampling strategy mitigates the risk of selecting trivial or overly simplistic negative samples, thereby reinforcing the model’s capacity to differentiate between nuanced reasoning paths and effectively learn from more challenging and contextually relevant negative samples.
+Generalized proximal policy optimization.
+Compared with PPO that optimizes the model by maximizing expected rewards through constrained policy updates, GRPO tang2024generalized extends PPO by incorporating additional objectives or multi-task learning capabilities, making it suited for financial reasoning tasks where accuracy and reasoning completeness, long-context reasoning abilities are crucial for effectively handling complex financial scenarios.
+Therefore, to comprehensively improve LLMs’ reasoning abilities in financial reasoning tasks, we propose a novel multi-faceted reward structure in GRPO that systematically evaluates four critical dimensions.
+Compared with prior work that predominantly targets accuracy and format chen2024huatuogpt ; finr1 ; zhu2025dianjin , our framework introduces two additional components: reasoning logic and length rewards.
+By integrating these dimensions, we aim to mitigate shortcut learning, reinforce logical consistency, and incentivize effective reasoning over long contexts—a crucial capability given the inherently lengthy and complex nature of financial documents.
+Formally, the reward function for GRPO is designed as follows:
+$r(x,y,y^{*})=\alpha_{1}\cdot r_{acc}+\alpha_{2}\cdot r_{logic}+\alpha_{3}\cdot r_{format}+\alpha_{4}\cdot r_{length}\cdot r_{acc}$ (3)
+where:
+  • $r_{acc}$: Accuracy reward, based on the verifier’s assessment of the final response correctness.
+  • $r_{logic}$: Logical consistency reward, evaluating the correctness of the reasoning path.
+  • $r_{format}$: Assessing if the generated response follows the designed format.
+  • $r_{length}$: To encourage the model’s ability to reason over lengthy contexts—a critical requirement in financial analysis given the extensive nature of many financial documents—we provide an additional reward when the answer is correct and the context length exceeds 8,192 tokens.
+If the accuracy reward is zero, the length reward is also set to zero.
+To effectively evaluate the generated response, we also leverage LLMs (GPT-4o) as judge.
+Following the reward score design in finr1 ; chen2024huatuogpt , we first assesses the accuracy of the generated response, signing a reward of 1 if the answer is deemed correct.
+Similarly, if the logical reasoning is validated as correct, a reward of 1 is also given.
+For format and length evaluation, we employ rule-based methods to assign corresponding rewards, the format reward is set as 0.1 and the length reward is set as 1.
+4 FinReason Evaluation Benchmark
+In this section, we present FinReason, the first comprehensive benchmark listed in Table 2 designed to systematically evaluate the financial reasoning capabilities of LLMs.
+4.1 Tasks
+Financial reasoning tasks present distinct challenges beyond general reasoning, as they require not only understanding the structure and semantic content of financial documents and tables but also accurately interpreting and executing complex equations.
+To this end, we introduce a unified benchmark that systematically integrates reasoning over (1) financial texts with tables, (2) long-context reasoning, (3) multi-table analysis, and (4) interpretation and execution of financial equations, establishing a comprehensive evaluation framework to assess and elevate LLM capabilities in financial reasoning tasks.
+Four datasets are selected to evaluate LLMs across these key aspects of financial reasoning:
+(1) FinQA chen2021finqa serves as the foundational benchmark for quantitative reasoning over structured tables and basic financial texts.
+It assesses the model’s ability to extract and synthesize financial information related to core financial concepts such as revenue, profit margins, and financial ratios.
+(2) Simplong zhao2024docmath targets long-context reasoning within single-table scenarios.
+It assesses the model’s capacity to perform numerical reasoning based on structured data while maintaining context throughout extended financial texts.
+(3) Complong zhao2024docmath extends Simplong by introducing multi-table scenarios, challenging the model to integrate information across multiple tables.
+This dataset tests the model’s ability to handle more complex financial reporting contexts, where financial data is fragmented and requires cross-table reasoning to draw accurate conclusions.
+(4) XBRL-Math wang2025finnlp emphasizes equation-based reasoning, requiring the model to accurately interpret and execute financial formulas derived from structured data in XBRL filings.
+This dataset assesses the model’s ability to associate financial terms with corresponding numerical data and perform structured calculations.
+Table 2: Overview of the datasets evaluated in the study.
+Dataset Size Data Types Data Source Avg. Token Metric License
+FinQA chen2021finqa 1,100 Tables and Texts Financial reports 1,128 Acc MIT License
+DM-Simplong zhao2024docmath 100 Tables and Texts Financial reports 4,330 Acc MIT License
+DM-Complong zhao2024docmath 300 Tables and Texts Financial reports 39,983 Acc MIT License
+XBRL-Math wang2025finnlp 90 Texts and Equations XBRL 397 Acc CDLA-Permissive 2.0
 4.2 Evaluated Models
-[…]
-The evaluated models are summarized as follows: GPT-4o Hurst et al.
-(2024): OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning.
-GPT-o1 Jaech et al.
-(2024): A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability.
-GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding.
-DeepSeek-V3 Liu et al.
-(2024): A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning.
-DeepSeek-R1 Guo et al.
-(2025): A first-generation reasoning model with competitive performance on math, code, and reasoning tasks.
-Qwen2.5-72B-Instruct Yang et al.
-(2024): A strong multilingual model with advanced instruction tuning.
-Qwen2.5-72B-Instruct-Math: A math-specialized variant fine-tuned for complex problem solving.
-LLaMA-series Dubey et al.
-(2024): Widely adopted open-source models (8B–70B) with strong instruction-tuned variants used in our experiments.
-DeepSeek-R1-Distill Guo et al.
-(2025): Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency.
-Other reasoning-enhanced models: Includes S1-32BMuennighoff et al.
-(2025), Limo Ye et al.
-(2025), and FinR1-7B Liu et al.
-(2025b), incorporating reasoning strategies such as financial CoT construction and task-specific fine-tuning.
-Model Name Parameters Enhanced for Reasoning Tasks Context Window Size Close/Open Source Reasoning Enhanced Training Data
-GPT-4o Unknown No 128k Closed -
-GPT-o1 Unknown Yes 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
-GPT-o3-mini Unknown Yes 128k Closed Public and proprietary data
-DeepSeek-V3 671B No 128k Open -
-DeepSeek-R1 671B Yes 128k Open Cold-start data generation, post-processing data
-GPT-4.5 Unknown No 128k Closed -
-Qwen2.5-72B-Instruct 72B No 128k Open -
-Qwen2.5-72B-Instruct-Math 72B Yes 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
-DeepSeek-R1-Distill-Llama-70B 70B Yes 128k Open Distilled from R1
-Llama3-70B-Instruct 70B No 8k Open -
-Llama3.1-70B-Instruct 70B No 128k Open -
-Llama3.3-70B-Instruct 70B No 128k Open -
-Qwen2.5-32B-Instruct 32B No 128k Open -
-DeepSeek-R1-Distill-Qwen-32B 32B Yes 128k Open Distilled from R1
-Limo 32B Yes 128k Open Public and proprietary data with Qwen filteration
-S1-32B 32B Yes 128k Open Public and proprietary data + original data, filtered by quality, difficulty, diversity
-Qwen/QwQ-32B 32B Yes 128k Open -
-Qwen2.5-14B-Instruct 14B No 128k Open -
-DeepSeek-R1-Distill-Qwen-14B 14B Yes 128k Open Distilled from R1
-DeepSeek-R1-Distill-Llama-8B 8B Yes 128k Open Distilled from R1
-Llama3-8B-Instruct 8B No 8k Open -
-Llama3.1-8B-Instruct 8B No 128k Open -
-Qwen2.5-7B-Instruct 7B No 128k Open -
-FinR1-7B 7B Yes 128k Open 60k CoT data for financial tasks
-Table 2: Summary of evaluated large language models, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies.
-MCTS refers to Monte Carlo Tree Search.
-CoT refers to chain-of-thought reasoning.
-TIR means tool-integrated reasoning.
+To investigate how existing LLMs, especially reasoning-enhanced LLMs, perform in financial tasks, we select a diverse set of 29 LLMs from the GPT, LLaMA, DeepSeek, and Qwen families, along with other reasoning models including both general domain and financial specific models.
+Our selection covers models of various scales, from smaller models (8B) to large-scale models (70B), and even ultra-large models like DeepSeek-R1 (671B).
+We also include advanced closed-source models such as GPT-4o and GPT-o1 as strong baselines.
+More details of the models can be found in Appendix E.
 4.3 Evaluation Settings
 In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question.
-Following prior work Zhao et al.
-(2024a), we use a generic prompt (see appendix D for more details) to instruct models to answer based on the given context.
-Models must extract relevant information and perform mathematical reasoning using numbers and concepts from the context.
-While outputs are typically in mathematical form, variations in formats (e.g., percentages, rounding) may cause differences.
-Reasoning-focused models like DeepSeek-R1 and GPT-o1 often generate step-by-step reasoning.
-Inspired by DocMath Zhao et al.
-(2024a), we adopt an LLM-as-judge evaluation approach: GPT-3.5-Turbo extracts final answers, followed by a mathematical comparison to assess correctness.
-For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series), we run evaluations via APIs from TogetherAI [^2] https://www.together.ai and OpenAI [^3] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06.
-For models up to 72B parameters, we use the VLLM framework on 8 A100 GPUs (80GB each), with a max generation length of 512 tokens.
-2.3.1 Reasoning Data Curation
-To enhance domain-specific reasoning capabilities in LLMs, we curate a high-quality training dataset by integrating diverse financial reasoning benchmarks.
-Our dataset combines a wide range of reasoning tasks—covering numerical computation, table-text integration, cross-table correlation, and long-context understanding—reflecting the complexity of real-world financial documents.
-Specifically, we incorporate data from several well-established benchmarks, including FinQA Chen et al.
-(2022), DocFinQA Reddy et al.
-(2024b), TAT-QA Zhu et al.
-(2021), DocMath-Eval Zhao et al.
-(2024b), and BizBench-QA Krumdick et al.
-(2024).
-These datasets collectively provide rich and diverse coverage across task types and difficulty levels, supporting the development of robust financial reasoning models.
-FinQA provides QA pairs combining textual and tabular financial data; DocFinQA extends this to long-context documents; TAT-QA focuses on complex numerical reasoning; DocMath-Eval (CompLong and SimpShort subsets) targets reasoning over multi-table and single-table scenarios; and BizBench-QA enriches the corpus with code-based financial reasoning tasks.
-Please see Appendix B for detailed descriptions of each dataset.
-2.4 Data Collection and Reasoning Path Construction
-To ensure a balanced dataset of both short and long context samples, we collect all data from the FinQA train set, TATQA, DocMath-CompLong, and DocMath-SimpShort.
-Additionally, we sample 1,000 examples each from BizBench-QA and DocFinQA to balance short and long reasoning tasks.
-In total, we gather 9,851 financial QA samples for reasoning path construction.
-Inspired by Ye et al.
-(2025); Muennighoff et al.
-(2025), which suggest that simple QA pairs may not effectively improve reasoning capabilities, we apply a filtering strategy to retain more challenging examples.
-Specifically, we use Qwen-14B-Instruct to generate answers for each question and employ GPT-3.5-Turbo as a verifier.
-QA pairs correctly answered by Qwen-14B are removed, resulting in a more difficult training subset of 3,472 QA pairs for optimizing more advanced models.
-2.5 Reasoning Path Generation and Refinement
-For reasoning path generation, we use GPT-4o, chosen for its strong performance across datasets (Table 3) and its stability in generating high-quality outputs.
-Following the CoT refinement framework from HuatuoGPT-o1 Chen et al.
-(2024), we employ a verifier-based iterative process to refine the reasoning paths generated by GPT-4o.
-If a reasoning path is rejected by the verifier, refinement is guided by one of the following strategies: (1) Backtracking: Revisiting previous steps to correct logical inconsistencies.
-(2) Exploring New Paths: Generating alternative reasoning routes when the current path fails.
-(3) Verification: Cross-checking intermediate steps to ensure logical soundness.
-(4) Correction: Making targeted adjustments to rectify specific errors.
-The CoT refinement process begins with an initial prompt to generate a reasoning path and final answer.
-If the verifier identifies the answer as incorrect, one of the above strategies is applied iteratively until a correct and logically consistent answer is produced.
-2.5.1 Model Training
-We develop our financial reasoning models based on curated reasoning paths, focusing on smaller-sized models for efficient real-world deployment.
-We select LLaMA-3.1-8B-Instruct and Qwen2.5-14B-Instruct as backbone models, given their strong performance among small and mid-sized LLMs.
-We adopt a two-stage training approach.
-In the first stage, we perform Supervised Fine-Tuning (SFT), using questions as input and reasoning paths with final answers as output.
-For the 8B model, we use only FinQA reasoning paths to assess the impact of domain-specific fine-tuning.
-For the 14B model, we use the full set of sampled and filtered data to leverage its greater capacity for diverse and complex reasoning.
-In the second stage, we apply RL using a verifier-based sparse reward mechanism and optimize with Proximal Policy Optimization (PPO), following Chen et al.
-(2024).
-A reward model evaluates the correctness and coherence of the reasoning path.
-If errors are found, the model iteratively refines its outputs through backtracking, alternative path exploration, verification, and correction, improving logical consistency and equation-solving accuracy.
-During SFT, models are trained for three epochs with a learning rate of 5e-6 and a maximum sequence length of 8192 tokens.
-LoRA is applied for efficient parameter updates.
-In the RL stage, training continues for three epochs with a reduced learning rate of 5e-7, guided by the reward model to refine reasoning quality via PPO.
+Following prior work wei2022chain , we use a generic prompt (see appendix G for more details) to instruct models to answer the question.
+While answers to financial tasks are often expressed in mathematical forms, variations in representation, such as differences in percentages, rounding conventions, or numeric formatting, can lead to discrepancies in surface-level comparisons.
+To address this, we opt out exact match metrics and instead adopt an LLM-as-judge evaluation paradigm.
+Specifically, we follow the answer extraction methodology proposed by chen2023theoremqa in TheoremQA, extracting the numerical results from model outputs.
+The extracted answers are then compared against the ground truth to assess correctness, allowing for flexible yet rigorous evaluation of generated responses.
+Please see appendix D.2 for more details of the evaluation setting.
 5 Results
-Table 3 presents the performance of various LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong.
-Our models, Fino1-8B and Fino1-14B, demonstrate strong and consistent results across tasks, outperforming several larger and closed-source models.
-Fino1-14B achieves an average score of 60.25, surpassing many larger-scale models such as LLaMA-3.3-70B-Instruct (56.04), Qwen2.5-72B-Instruct (53.71), GPT-o1-preview (54.05), GPT-o3-mini (57.89), and even GPT-4.5 (60.43).
-It also achieves the best performance among all models on two benchmarks—DM-Simplong (60.00) and XBRL-Math (86.67).
-These results demonstrate that high-quality, domain-specific reasoning training can enable smaller models to rival—and even surpass—the performance of significantly larger and proprietary LLMs.
-Fino1-8B also performs competitively with an average score of 50.77, surpassing other 8B models and even some 14B and 32B models.
-Both models excel on XBRL-Math, where Fino1-14B achieves the top score of 86.67.
-Moreover, Fino1-14B and Fino1-8B significantly outperform FinR1-7B, a financial reasoning-enhanced model trained on a much larger 60K task-aligned dataset, demonstrating that larger data volume does not necessarily lead to better performance, and that carefully curated, high-quality reasoning data—even in smaller quantities—can yield superior results.
+Next, we present the results of these LLMs on FinReason and discuss our main findings through the results for the research questions listed in section  1.
+Table 3: Overall performance of different LLMs on four financial reasoning datasets.
 Models FinQA DM-Simplong XBRL-Math DM-Complong Average
@@
 GPT-4.5 68.94 59.00 74.44 39.33 60.43
-Meta-Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
+Llama-4-Scout 70.45 52.00 88.89 0.67 53.00
+Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
 Llama-3.1-70B-Instruct 63.18 48.00 63.33 34.33 52.21
@@
 Qwen2.5-32B-Instruct 73.11 56.00 65.56 30.00 56.17
-Qwen/QwQ-32B 61.22 46.00 84.44 20.00 52.92
+Qwen3-32B 64.15 51.00 85.56 26.00 56.68
+QwQ-32B 61.22 46.00 84.44 20.00 52.92
 DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 24.67 57.40
@@
 Qwen2.5-14B-Instruct 67.44 59.00 57.78 26.67 52.72
+Qwen3-14B 64.33 49.00 86.67 24.00 56.00
 DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 21.00 53.18
 DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 15.67 43.94
-Meta-Llama-3-8B-Instruct 41.97 29.00 48.89 6.00 31.47
+Llama-3-8B-Instruct 41.97 29.00 48.89 6.00 31.47
 Llama-3.1-8B-Instruct 54.13 34.00 62.22 14.30 41.16
 Qwen2.5-7B-Instruct 55.37 41.00 42.22 17.67 39.07
+Qwen3-8B 62.11 46.00 83.33 17.67 52.28
 FinR1-7B 58.74 37.00 30.00 13.67 34.85
-Fino1-8B 60.87 40.00 82.22 20.00 50.77
-Fino1-14B 70.01 60.00 86.67 24.33 60.25
-Table 3: Performance of different LLMs on four financial reasoning datasets.
-Beyond overall performance, the comparison across models provides additional insights that can guide the development of LLMs for financial applications:
-(1) General reasoning enhancements do not necessarily improve financial task performance.
-Models like GPT-o1 and DeepSeek-R1, though strong in general reasoning, underperform on financial benchmarks, highlighting the importance of domain-specific knowledge and structured numerical reasoning.
-Moreover, reasoning enhancement strategies based on general-domain data, as seen in S1-32B and Limo, often fail to improve and sometimes even degrade performance compared to their backbone models (e.g., Qwen2.5-32B).
-These results underscore that high-quality, domain-aligned reasoning paths are more effective than general reasoning augmentation, and tailored financial training is crucial for optimal model performance.
-(2) Reasoning enhancements vary in effectiveness across financial tasks.
-Models like GPT-o1 perform worse than GPT-4o on FinQA (49.07 vs.
-72.49) and DM-Simplong (56.00 vs.
-60.00).
-Qwen2.5-Math improves in XBRL-Math (67.78 $\rightarrow$ 83.33) but drops sharply in FinQA (69.74 $\rightarrow$ 42.00) and DM-Simplong (59.00 $\rightarrow$ 42.00).
-DeepSeek-R1-Distill-Llama-8B also excels in XBRL-Math (81.11) but shows limited gains elsewhere.
-These results suggest that strategies like CoT or math fine-tuning may offer benefits for structured tasks but often fail to generalize effectively to broader financial reasoning.
-Compared with Qwen2.5-72B-Instruct and its math enhanced version, we can see that while TIR and synthetic building appears to help in equation-heavy tasks like XBRL-Math, its impact on other types of financial tasks is less clear and may be influenced by multiple factors such as task structure, data modality, or domain complexity.
-(3) Larger model size offers limited benefits in financial reasoning.
-While larger models typically excel in general NLP tasks, their gains in financial tasks are not proportional.
-For example, DeepSeek-70B (66.73 FinQA, 53 DM-Simplong, 86.67 XBRL-Math) performs similarly to the smaller DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite having less than half the parameters.
-Across models like Qwen2.5-72B, DeepSeek-R1-Distill-Llama-70B, and GPT-4o, performance tends to plateau around 70B.
-This may reflect the structured nature of tasks like XBRL-Math, where arithmetic precision outweighs complex language reasoning, enabling smaller models to compete effectively.
-(4) Improved pre-training data and strategies in general domains also benefit financial tasks.
-Llama 3.1-8B-Instruct shows a notable performance boost over Llama 3-8B-Instruct (avg.
-50.12 vs.
-39.95), with gains in XBRL-Math (62.22 vs.
-48.89) and FinQA (54.13 vs.
-41.97).
-Similarly, Llama 3.1-70B-Instruct outperforms Llama 3-70B (58.17 vs.
-52.20), especially in DM-Simplong (48.00 vs.
-41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05), driven by improvements in FinQA (68.15) and XBRL-Math (70.00).
-These gains stem from larger, more diverse training corpora (e.g., 15T+ tokens in Llama 3.1) and enhanced post-training techniques like online preference optimization.
-Such strategies improve general reasoning while also boosting performance in domain-specific tasks like financial reasoning.
-(5) Existing reasoning models struggle with long contexts and cross-table correlations.
-Financial tasks often require integrating information from lengthy reports, structured tables, and equations, yet current models show clear limitations.
-For example, GPT-4o performs well on FinQA (72.49) but drops on DM-Simplong (60.00), indicating challenges in combining textual and symbolic reasoning.
-Models like DeepSeek-R1 and DeepSeek-R1-Distill-Qwen-32B excel in XBRL-Math but fail to generalize.
-Performance on DM-Complong remains low across models (e.g., LLaMA-3.3-70B: 32.00).
-Smaller models like DeepSeek-R1-Distill-Llama-8B also perform poorly with 33.00 for DM-Simplong.
-These results highlight the need for improved long-context modeling and structured data integration in financial reasoning.
-(6) Financial reasoning enhancement improves performance across all tasks.
-Compared to LLaMA 3.1-8B-Instruct (41.16), DeepSeek-R1-Distill-Llama-8B shows a modest gain (43.94).
-In contrast, our Fino1-8B, trained on GPT-4o-distilled reasoning paths with backtracking strategies, achieves a substantial improvement, reaching an average score of 50.77, and consistently outperforms baselines on FinQA (60.87 vs.
-54.13), DM-Simplong (40.00 vs.
-34.00), and XBRL-Math (82.22 vs.
-62.22).
-Fino1-14B further boosts performance, achieving an average score of 60.25, outperforming other 14B models (e.g., Qwen2.5-14B: 52.72) and even rivaling state-of-the-art models like DeepSeek-R1 (60.87) and GPT-4o (61.01).
-Notably, although Fino1-8B and Fino1-14B were trained with limited financial CoTs, both models show improvements across all tasks, highlighting the effectiveness of high-quality reasoning paths and structured training strategies such as backtracking, path exploration, and verification.
-Overall, these insights highlight that financial reasoning poses unique challenges distinct from general NLP and mathematical reasoning.
-The performance variability across tasks underscores the need for specialized training strategies and architectural improvements to better support financial comprehension, structured data processing, and multi-step numerical reasoning.
+Dianjin-R1-7B 60.20 35.00 83.33 14.67 48.30
+Fino1-8B 73.03 56.00 84.44 26.33 59.95
+Fino1-14B 74.18 55.00 87.78 27.33 61.07
+Table 3 presents the performance of evaluated LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong.
+Our Fin-o1-14B achieves a score of 61.07, ranking second only to Deepseek-V3 despite having substantially fewer parameters.
+Notably, Fin-o1-14B surpasses larger-scale models like LLaMA-3.3-70B-Instruct, Qwen2.5-72B-Instruct, even GPT-o1-preview, GPT-o3-mini, and GPT-4.5, underscoring its superior financial reasoning capabilities.
+Fin-o1-8B also performs competitively, attaining an average score of 58.58, exceeding most models with 70B parameters or fewer.
+Both Fin-o1-8B and Fin-o1-14B excel in FinQA, DM-Simplong, and XBRL-Math, demonstrating the impact of targeted domain adaptation in financial reasoning tasks.
+These gains also confirm the effectiveness of our three-stage FinCoT construction and the two-phase SFT + RL training strategy.
+General reasoning LLMs fail in financial reasoning.
+While general reasoning models exhibit significant improvements in general mathematical reasoning tasks, they struggle to effectively transfer these gains to financial reasoning tasks, which demand specialized knowledge and reasoning over financial documents and tables.
+For instance, the performance of QwQ-32B and Limo declined notably compared to Qwen2.5-32B-Instruct, dropping from 56.17% to 52.92% and 46.22%, respectively.
+Similarly, although S1-32B demonstrated overall performance gains, its effectiveness in FinQA, Simplong and Complong decreased.
+These results highlight the persistent gap between general and financial reasoning due to the complexities of interpreting financial tables and structures.
+Incomprehensive reasoning data limits consistent improvements.
+While our model demonstrates consistent improvements over the backbone model Qwen3 across all datasets, other financial models do not exhibit such stable improvements.
+For instance, while Fin-R1 and Dianjin-R1-7B achieved higher scores in FinQA, their performance in long-context reasoning tasks dropped significantly compared with their backbone model Qwen-2.5-7B-Instruct.
+On Simplong, the backbone model scored 41%, but Fin-R1 and Dianjin-R1-7B fell to 37% and 35%, respectively.
+A similar decline is observed on Complong, indicating the challenges in maintaining long context reasoning performance.
+These declines are attributable to CoT datasets used in these works, which focused on shorter contexts and knowledge-based QA rather than reasoning over contexts, limiting the models’ exposure to reasoning over extended financial contexts.
+Table 4: Ablation study of Fin-o1-8B with different SFT training data and RL methods.
+Models FinQA DM-Simplong XBRL-Math DM-Complong Average
+Backbone model
+Qwen3-8B 62.11 46.00 83.33 17.67 52.28
+SFT model
+Fin-o1-8B (SFT with original QA pair) 66.55 50.00 74.44 16.67 51.92
+Fin-o1-8B (SFT with all CoT data) 71.32 51.00 81.11 21.67 56.28
+Fin-o1-8B (SFT with filtered CoT) 71.78 57.00 82.22 23.33 58.58
+SFT + RL model
+Fin-o1-8B (SFT + PPO) 72.05 54.00 83.33 23.00 58.10
+Fin-o1-8B (SFT + DPO) 69.74 51.00 72.22 25.00 54.49
+Fin-o1-8B (SFT + GRPO) 73.03 56.00 84.44 26.33 59.95
+Table 4 presents additional results evaluating the effectiveness of various RL methods in financial reasoning tasks and further verifying the effectiveness of FinCoT.
+Based on the results, we find that:
+GRPO Outperforms Other RL Methods in Financial Reasoning.
+GRPO achieves the most substantial gains, driven by its multi-faceted reward function that emphasizes correct reasoning and long-context comprehension.
+Notably, in DM-Complong, where other models like Fin-R1 and S1-32B falter, the SFT+GRPO model reaches 26.33% with just 8B parameters—a 9% improvement over its backbone model.
+In contrast, PPO shows gains in FinQA and XBRL-Math but declines in Simplong and Complong, indicating weaker long-context reasoning.
+For DPO, while the performance in DM-Complong improves, performance in simpler tasks declines.
+This may stem from the design of positive-negative sample pairs, where the negative path is curated from the last-mistake path and the positive path is generated by allowing the model to take one additional reasoning step.
+This approach may inadvertently encourage excessive reasoning, benefiting complex tasks but potentially causing overthinking in simpler tasks.
+We further compare the impact of fine-tuning the backbone model using original QA pairs without CoT reasoning and with unfiltered CoT data.
+Training without CoT data slightly improves basic financial reasoning tasks like FinQA, but leads to performance drops in XBRL-Math and Complong.
+This decline is particularly pronounced in XBRL-Math, where the lack of step-by-step reasoning impedes the model’s ability to handle financial equations, hindering its generalization to equation-heavy tasks.
+In contrast, using the unfiltered FinCoT dataset demonstrates the effectiveness of exposing the model to more complex reasoning paths, underscoring the importance of reasoning-focused training for financial tasks.
 6 Conclusion
-This study presents a comprehensive evaluation of 24 general-purpose and reasoning-augmented LLMs on financial reasoning tasks, encompassing financial text understanding, numerical computation, and structured data interpretation.
-Our findings show that existing models, despite strong performance on general reasoning benchmarks, often fail to generalize to financial domains—particularly in handling long contexts, multi-table reasoning, and domain-specific terminology.
-To bridge this gap, we curated a high-quality, diverse financial reasoning dataset and introduced Fino1, a suite of domain-adapted models trained with GPT-4o-generated reasoning paths using chain-of-thought fine-tuning and reinforcement learning.
-Despite being trained on limited data, Fino1 models outperform several larger and advanced LLMs, including GPT-o1, GPT-o3-mini, GPT-4.5, and DeepSeek models (V3 and R1).
-These results highlight the effectiveness of domain-specific reasoning strategies and demonstrate the practical value of tailored model adaptation for real-world financial applications.
+In this work, we address the critical yet underexplored challenge of financial reasoning for LLMs.
+We present three key contributions to bridge existing gaps: (1) the development of FinCoT, the first publicly available CoT dataset specifically curated for financial reasoning tasks; (2) the introduction of Fin-o1, the first open-source financial reasoning model trained through a two-phase framework combining supervised fine-tuning and RL methods; and (3) the establishment of FinReason, a comprehensive benchmark designed to systematically evaluate financial reasoning across essential components such as multi-table analysis, long-context reasoning, and equation-based tasks.
+Our extensive evaluations demonstrate that general reasoning models struggle to adapt to financial contexts, revealing significant overall performance drops.
+In contrast, our Fin-o1 models, trained on FinCoT with GRPO, achieve consistent gains across diverse financial reasoning tasks, highlighting the critical role of domain-specific CoT supervision and structured RL optimization.
+These findings underscore the need for targeted datasets, specialized models, and advanced training strategies to effectively address the complexities of financial reasoning, setting a new standard for future research in this domain.
 Appendix A Related Work
 Our work builds upon three main research directions: the development of general-purpose and reasoning-enhanced language models, applications of LLMs in finance, and studies on reasoning capabilities in financial tasks.
-A.1 General-Purpose and Reasoning-Enhanced LLMs
+A.1 General-Purpose and Reasoning LLMs
 Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning.
-Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting Wei et al.
-(2022), self-reflection Zhang et al.
-(2023), and reinforcement learning from human feedback Ouyang et al.
-(2022).
-These models have shown superior performance in mathematical reasoning Lewkowycz et al.
-(2022), symbolic manipulation Drori et al.
-(2023), and logical inference tasks Chen et al.
-(2023).
+Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting [35], self-reflection [44], and reinforcement learning from human feedback [23].
+These models have shown superior performance in mathematical reasoning [16], symbolic manipulation [7], and logical inference tasks [2].
 A.2 LLMs in Financial Applications
-The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis Yang et al.
-(2020), market sentiment prediction Yang et al.
-(2023), and automated trading strategies Kou et al.
-(2024).
-Recent work has focused on developing finance-specific models like BloombergGPT Wu et al.
-(2023), FinGPT Yang et al.
-(2023) and PIXIU Xie et al.
-(2023), OpenFinLLMs Xie et al.
-(2024b), which are pretrained on financial corpora to better understand domain-specific terminology and concepts.
+The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis [42], market sentiment prediction [41], and automated trading strategies [13].
+Recent work has focused on developing finance-specific models like BloombergGPT [36], FinGPT [41] and PIXIU [38], OpenFinLLMs [39], Plutus[24], which are pretrained on financial corpora to better understand domain-specific terminology and concepts.
 These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
-A.3 Reasoning Capabilities in Financial Tasks
+A.3 Financial Reasoning
 Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference.
-Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al.
-(2021), mathematical problem-solving in financial contexts Srivastava et al.
-(2024), and structured data interpretation Liu et al.
-(2025a).
+Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents [4], mathematical problem-solving in financial contexts [29], and structured data interpretation [18].
 These works highlight the importance of both domain expertise and reasoning abilities in financial applications.
 Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
-Appendix B Details Data Source
+Additionally, we focus on benchmarking these models across diverse financial datasets, emphasizing multi-table reasoning, long-context comprehension, and quantitative analysis.
+This comprehensive evaluation not only sheds light on current model limitations but also informs future research directions for developing more robust financial reasoning models.
+Appendix B Details of Data Source
+B.1 Data source examples for other works
+This section present some examples of the data from other models. In convince of reviewing, we have translated all the original context into English if they are not English.
+Finance-Instruct Explain tradeoffs between fiscal and monetary policy as tools in a nations economic toolkit.
+Provide examples of past instances when each were utilized, the economic conditions that led to them being deployed, their intended effects, and an evaluation of their relative efficacy and consequences.
+CFLUE Assume you are a financial industry expert.
+Please answer the following question.
+Note: The question is a single-choice question.
+Only return the most appropriate option.
+If multiple options are suitable, only return the most accurate one.
+Among the following options, which financial institution is not permitted to participate in the interbank lending market?
+A.
+Clearing Center B.
+Insurance Company C.
+Securities Company D.
+Fund Management Company
+CFLUE Is the sentiment of this text positive, negative, or neutral? Feeling exhausted from averaging down.
+Finance-IQ When considering the pricing of life insurance products, several key factors come into play, including market forces, business operating costs, and government regulations.
+Among these factors, which one primarily determines the minimum limit of life insurance product prices?
+A.
+Market forces and business operating costs B.
+Business operating costs C.
+Government regulations D.
+Market forces
+B.2 Data source for FinCoT
 We provide here detailed descriptions of the datasets used to construct our training and evaluation corpus for financial reasoning enhancement.
-  • FinQA Chen et al.
-(2022): FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables.
+  • FinQA [5]: FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables.
 It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
-  • DocFinQA Reddy et al.
-(2024b): An extension of FinQA designed for long-context reasoning.
+  • DocFinQA [27]: An extension of FinQA designed for long-context reasoning.
 It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words.
 This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
-  • TAT-QA Zhu et al.
-(2021): A large-scale financial QA dataset created from real financial reports.
+  • TAT-QA [47]: A large-scale financial QA dataset created from real financial reports.
 It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison.
 TAT-QA reflects real-world challenges in financial document interpretation.
-  • DocMath-Eval Zhao et al. (2024b): This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
-    • DocMath-CompLong: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
-    • DocMath-SimpShort: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
-  • BizBench-QA Krumdick et al.
-(2024): Contains 14,400 financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode.
+  • ConvFinQA [6]: A financial question answering dataset designed to handle multi-turn conversations involving financial tables and text.
+ConvFinQA builds upon FinQA by incorporating follow-up questions and context retention across multiple turns, simulating more realistic financial analysis scenarios.
+It provides expert-annotated sub-questions that decompose complex financial questions into granular steps, facilitating the generation of structured reasoning paths and promoting systematic financial reasoning.
+  • DocMath-Eval [45]: This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
+    • DocMath-CompShort[45]: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
+    • DocMath-SimpShort[45]: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
+  • Econ_Logic [25]: It consists of financial logic reasoning tasks that help the model to further understand and have more insights of financial terms and knowledge.
+  • BizBench-QA [14]: Contains financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode.
 These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
+FinQA Please answer the given financial question based on the context.
+Context: measurement point december 31 the priceline group nasdaq composite index s&p 500 rdg internet composite .
+|measurement pointdecember 31|the priceline group inc .|nasdaqcomposite index|s&p 500index|rdg internetcomposite| |2010|100.00|100.00|100.00|100.00| |2011|117.06|100.53|102.11|102.11| |2012|155.27|116.92|118.45|122.23| |2013|290.93|166.19|156.82|199.42| |2014|285.37|188.78|178.29|195.42| |2015|319.10|199.95|180.75|267.25|.
+Question: what was the percent of the growth of the the priceline group inc .
+from 2014 to 2015 Answer:
+Appendix C Example of CoT data curation
+Here is an example of applying the guideline and iterative refinement process to generate a CoT reasoning path based on the given context and question.
+Figure 3: Example of iterative refinement of the CoT process with a guidance.
+BizBench From the results reflected in the following pro forma information because of normal production declines, changes in commodity prices, future acquisitions and divestitures, future development and exploration activities and other factors.
+| | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | | | (In millions, except per share amounts) | | Six Months Ended June 30, 2021 | | Pro forma revenue | | \$ | 1,802 | | | | | Pro forma net income | | 71 | | | | | | | | | | | | | | | | | | | | | | | | | | | | | Pro forma basic earnings per share | | \$ | 0.09 | | | | | Pro forma diluted earnings per share | | \$ | 0.09 | | | | 3.
+Properties and Equipment, NetProperties and equipment, net are comprised of the following: | | | | | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | (In millions) | | June 30,2022 | | December 31,2021 | | Proved oil and gas properties | | \$ | 16,102 | | | \$ | 15,340 | | | Unproved oil and gas properties | | 5,292 | | | 5,316 | | | Gathering and pipeline systems | | 423 | | | 395 | | | Land, buildings and other equipment | | 144 | | | 140 | | | Finance lease right-of-use asset | | 24 | | | 20 | | | | | 21,985 | | | 21,211 | | | Accumulated depreciation, depletion and amortization | | (4,578) | | | (3,836) | | | | | \$ | 17,407 | | | \$ | 17,375 | | As of June 30, 2022, the Company had no projects with exploratory well costs capitalized for more than one year after drilling.
+| | | | | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | (In millions) | | June 30,2022 | | December 31,2021 | | 6.51% weighted-average private placement senior notes (1) | | \$ | 37 | | | \$ | 37 | | | 5.58% weighted-average private placement senior notes (2) | | 87 | | | 87 | | | 3.65% weighted-average private placement senior notes | | 825 | | | 825 | | | 4.375% senior notes due June 1, 2024 | | 750 | | | 750 | | | 3.90% senior notes due May 15, 2027 | | 750 | | | 750 | | | 4.375% senior notes due March 15, 2029 | | 500 | | | 500 | | | Revolving credit facility | | - | | | - | | | Net premium (discount) | | 164 | | | 185 | | | Unamortized debt issuance costs | | (8) | | | (9) | | | | | \$ | 3,105 | | | \$ | 3,125 | | (1) Includes \$37 million of current portion of long-term debt at June 30, 2022.(2) Includes \$87 million of current portion of long-term debt at June 30, 2022.At June 30, 2022, the Company was in compliance with all financial and other covenants for both its revolving credit facility and senior notes.
+Question: What is the pro forma net income?
+Appendix D Details of experiments
+D.1 Model training
+We list the details of the training process here.
+The SFT process is configured with a learning rate of 5e-6 and a weight decay of 0.1.
+Training is conducted for 3 epoch with a batch size of 2 per GPU and a gradient accumulation of 2 steps, effectively resulting in an effective batch size of 2.
+The gradient checkpointing is enabled to optimize memory usage.
+For RL, the training process is configured with a learning rate specified set to 5e-6.
+The gradient accumulation is set to 2 steps, and a gradient accumulation of 2 steps, effectively resulting in an effective batch size based on the configured gradient_accumulation_steps.
+The training is conducted using 3 GPUs, each handling a separate training process.
+And a total of 500 steps is used to train the model.
+D.2 Model evaluation
+For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series, Llama4), we run evaluations via APIs from TogetherAI [^4] https://www.together.ai and OpenAI [^5] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06.
+All open-sourced models up to 72B parameters were conducted with vLLM [15] framework on 8 A100 GPUs (80GB each), with a max generation length of 1,024 tokens and temperature setting of 0.
+Given the extreme length of some evaluated dataset (e.g., Dm-Complong with a average length of 40K tokens), we use a zero-shot setting, aim to evaluating LLMs’ capabilities for financial reasoning without any few-shot demonstrations.
+Appendix E Details of Evaluated Models
+To comprehensively evaluate the performance of LLMs in financial tasks, particularly those with reasoning enhancements, we select a diverse set of 29 LLMs.
+This selection includes both general-purpose, and finance-specific models, ranging from smaller models (8B) to large-scale models (70B), and extending to ultra-large models such as DeepSeek-R1 (671B), GPT-4o, and GPT-o1.
+The evaluated models are summarized as follows: GPT-4o [11]: OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning.
+GPT-o1 [12]: A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability.
+GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding.
+DeepSeek-V3 [17]: A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning.
+DeepSeek-R1 [10]: A first-generation reasoning model with competitive performance on math, code, and reasoning tasks.
+Qwen series [40]: A strong multilingual model with advanced instruction tuning.
+LLaMA-series [8]: Widely adopted open-source models (8B–70B) as well as the latest MOE models with strong instruction-tuned variants used in our experiments.
+DeepSeek-R1-Distill [10]: Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency.
+Financial-Specific Reasoning Models: This category includes models such as Fin-R1-7B [19] and Dianjin-R1-7B [48], both of which leverage financial reasoning datasets and employ fine-tuning and reinforcement learning techniques to optimize LLM performance in financial reasoning tasks.
+Other reasoning models: They include S1-32B [22], Limo [43].
+These two models are designed for general reasoning tasks, with limited but high quality COT data.
+More details of the evaluated models are presented in Table 5.
+It also highlights the diverse training strategies employed across the evaluated models.
+Reinforcement learning has been extensively utilized to enhance reasoning capabilities, with common methods including PPO, DPO, and GRPO.
+Among these, GRPO is the most widely adopted, as seen in models such as DeepSeek-R1, Qwen2.5-72B-Instruct, and FinR1-7B.
+This widespread use of GRPO underscores its effectiveness in refining reasoning abilities, particularly in complex, multi-step tasks involving financial documents.
+Table 5: Summary of evaluated LLMs, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies.
+MCTS refers to Monte Carlo Tree Search.
+CoT refers to chain-of-thought reasoning.
+TIR means tool-integrated reasoning.
+Model Name Parameters Training Window Size Close/Open Source Reasoning Training Data
+GPT-4o - - 128k Closed -
+GPT-o1 - RL with long CoT 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
+GPT-o3-mini - - 128k Closed Public and proprietary data
+DeepSeek-V3 671B GRPO 128k Open -
+DeepSeek-R1 671B GRPO 128k Open Cold-start data generation, post-processing data
+GPT-4.5 - - 128k Closed -
+Llama4-Scout 17B*16E SFT & DPO 128k Open -
+Qwen2.5-72B-Instruct 72B DPO & GRPO 128k Open -
+Qwen2.5-72B-Instruct-Math 72B DPO & GRPO 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
+DeepSeek-R1-Distill-Llama-70B 70B GRPO 128k Open Distilled from R1
+Llama3-70B-Instruct 70B SFT & DPO 8k Open -
+Llama3.1-70B-Instruct 70B SFT & DPO 128k Open -
+Llama3.3-70B-Instruct 70B SFT & DPO 128k Open -
+Qwen2.5-32B-Instruct 32B DPO & GRPO 128k Open -
+Qwen3-32B 32B RL with long CoT 128k Open -
+DeepSeek-R1-Distill-Qwen-32B 32B GRPO 128k Open Distilled from R1
+Limo 32B SFT 128k Open Public data with Qwen filteration, with CoT generated from Deepseek
+S1-32B 32B SFT 128k Open Public data and original data, filtered by quality, with CoT generated from Gemini
+Qwen/QwQ-32B 32B PPO 128k Open -
+Qwen2.5-14B-Instruct 14B DPO & GRPO 128k Open -
+Qwen3-14B 14B RL with long CoT 128k Open -
+DeepSeek-R1-Distill-Qwen-14B 14B GRPO 128k Open Distilled from R1
+DeepSeek-R1-Distill-Llama-8B 8B GRPO 128k Open Distilled from R1
+Llama3-8B-Instruct 8B SFT & DPO 8k Open -
+Llama3.1-8B-Instruct 8B SFT & DPO 128k Open -
+Qwen2.5-7B-Instruct 7B DPO & GRPO 128k Open -
+Qwen3-8B 8B - 128k Open -
+FinR1-7B 7B SFT & GRPO 128k Open 60k CoT data for financial tasks
+Dianjin-R1-7B 7B SFT & GRPO 128k Open 31k CoT data for financial tasks
 Figure 4: Error case 1.
 Figure 5: Error case 2.
-Appendix C Error Analysis
+Appendix F Error Analysis
 To further investigate why reasoning-enhanced models like DeepSeek-R1 and GPT-o1 underperform in financial tasks compared to general-purpose models such as GPT-4o and DeepSeek-V3, we conduct an error analysis using the FinQA dataset since these reasoning-enhanced models exhibit a significant performance drop on FinQA.
@@
 This highlights a fundamental gap in DeepSeek-R1’s financial reasoning capabilities, particularly in recognizing the accounting conventions used to report share-based compensation.
-Limitations
-Our study has several limitations that highlight areas for future improvement.
-First, our model scale is limited, as we only fine-tuned an 8B model (Fino1), while larger models (e.g., 70B) could potentially benefit more from reasoning enhancements.
-Second, our evaluation scope is restricted, covering only three financial reasoning tasks (FinQA, DM-Simplong, and XBRL-Math), which do not fully capture the breadth of financial NLP applications such as forecasting, financial sentiment analysis, and fraud detection.
-Third, our fine-tuning relies on a single dataset, FinQA, for reasoning path construction, limiting the model’s exposure to different financial reasoning patterns; incorporating additional datasets could improve generalization.
-Fourth, our reasoning path construction approach is simplified, as we generate paths using a single method (GPT-4o), whereas exploring multiple reasoning path generation strategies—such as ensemble approaches or human-annotated paths—could lead to more robust financial reasoning capabilities.
-C.1 Future Direction
-Developing effective reasoning models for financial tasks poses distinct challenges compared to general-domain reasoning.
-Our findings suggest three key directions for improvement: (1) Enhancing financial knowledge and terminology understanding through domain-specific pretraining using financial corpora, structured reports, and regulatory documents; (2) Improving multi-table reasoning and long-context comprehension by developing strategies that better infer logic across tables and optimize memory mechanisms for processing lengthy financial documents; and (3) Refining reasoning-enhanced strategies, as current approaches like GPT-o1’s chain-of-thought prompting and DeepSeek-R1’s self-reflection via reinforcement learning yield mixed results—highlighting the need for more tailored reasoning strategies to address both structured numerical and unstructured textual financial reasoning.
-Ethics Statement
-While our study demonstrates significant advancements in financial reasoning with LLMs, Fino1 remains a research prototype and is not yet suitable for real-world financial applications.
-Our model inherits the well-documented limitations of large language models, including hallucinations, sensitivity to input phrasing, and potential biases in financial data sources, which could affect reliability in critical financial contexts.
-Furthermore, Fino1 has not undergone rigorous testing for high-stakes financial decision-making, such as investment strategies, regulatory compliance, or risk assessment, where incorrect outputs could lead to severe consequences.
-Additionally, the model’s reliance on a single dataset (FinQA) and a single reasoning path construction method (GPT-4o) limits its robustness and adaptability to the complex and evolving nature of financial environments.
-As with all AI-driven financial models, careful human oversight remains essential, and future work must focus on enhancing factual consistency, mitigating biases, improving domain adaptation, and ensuring alignment with regulatory frameworks before such models can be considered for real-world financial deployment.
-Appendix D Prompt
+Appendix G Prompt
 Prompt Template for DocMath-CompLong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context.
@@
 OUTPUT:
+Limitations
+Our study has several limitations that highlight areas for future improvement.
+First, our model scale is limited, as we only fine-tuned an 8B model (Fino1), while larger models (e.g., 70B) could potentially benefit more from reasoning enhancements.
+Second, our evaluation scope is restricted, covering only three financial reasoning tasks (FinQA, DM-Simplong, and XBRL-Math), which do not fully capture the breadth of financial NLP applications such as forecasting, financial sentiment analysis, and fraud detection.
+Third, our fine-tuning relies on a single dataset, FinQA, for reasoning path construction, limiting the model’s exposure to different financial reasoning patterns; incorporating additional datasets could improve generalization.
+Fourth, our reasoning path construction approach is simplified, as we generate paths using a single method (GPT-4o), whereas exploring multiple reasoning path generation strategies—such as ensemble approaches or human-annotated paths—could lead to more robust financial reasoning capabilities.
+Future Direction
+Developing effective reasoning models for financial tasks poses distinct challenges compared to general-domain reasoning.
+Our findings suggest three key directions for improvement: (1) Enhancing financial knowledge and terminology understanding through domain-specific pretraining using financial corpora, structured reports, and regulatory documents; (2) Improving multi-table reasoning and long-context comprehension by developing strategies that better infer logic across tables and optimize memory mechanisms for processing lengthy financial documents; and (3) Refining reasoning-enhanced strategies, as current approaches like GPT-o1’s chain-of-thought prompting and DeepSeek-R1’s self-reflection via reinforcement learning yield mixed results—highlighting the need for more tailored reasoning strategies to address both structured numerical and unstructured textual financial reasoning.
+Ethics Statement
+While our study demonstrates significant advancements in financial reasoning with LLMs, Fino1 remains a research prototype and is not yet suitable for real-world financial applications.
+Our model inherits the well-documented limitations of large language models, including hallucinations, sensitivity to input phrasing, and potential biases in financial data sources, which could affect reliability in critical financial contexts.
+Furthermore, Fino1 has not undergone rigorous testing for high-stakes financial decision-making, such as investment strategies, regulatory compliance, or risk assessment, where incorrect outputs could lead to severe consequences.
+Additionally, the model’s reliance on a single dataset (FinQA) and a single reasoning path construction method (GPT-4o) limits its robustness and adaptability to the complex and evolving nature of financial environments.
+As with all AI-driven financial models, careful human oversight remains essential, and future work must focus on enhancing factual consistency, mitigating biases, improving domain adaptation, and ensuring alignment with regulatory frameworks before such models can be considered for real-world financial deployment.
+Potential impacts
+The proposed work presents both positive and negative societal implications.
+The development of advanced financial reasoning models has the potential to enhance financial analysis by providing more accurate predictions and robust risk assessments, thereby supporting informed decision-making in areas such as investment strategies and regulatory compliance.
+However, these models may also pose risks if misused, such as reinforcing existing biases in financial datasets or facilitating manipulative financial practices.
+To mitigate such risks, we incorporate safeguards that enforce controlled access to the models, implement guidelines for responsible use, and maintain transparency in model outputs.
+Additionally, we emphasize the importance of adhering to ethical guidelines throughout the development and deployment process to minimize unintended societal consequences.